### v2

# Statistical Test on Diffusion Model-based Generated Images by Selective Inference

Abstract
AI technology for generating images, such as diffusion models, has advanced rapidly. However, there is no established framework for quantifying the reliability of AI-generated images, which hinders their use in critical decision-making tasks, such as medical image diagnosis. In this study, we propose a method to quantify the reliability of decision-making tasks that rely on images produced by diffusion models within a statistical testing framework. The core concept of our statistical test involves using a selective inference framework, in which the statistical test is conducted under the condition that the images are produced by a trained diffusion model. As a case study, we study a diffusion model-based anomaly detection task for medical images. With our approach, the statistical significance of medical image diagnostic outcomes can be quantified in terms of a $p$-value, enabling decision-making with a controlled error rate. We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
1 Introduction
Advances in image generation AI, such as diffusion models, have been remarkable (Song and Ermon, 2019). They can generate highly realistic and detailed images, which leads to innovations in various tasks across various fields. For example, image generation AI significantly enhances medical image diagnosis by improving accuracy and efficiency. It can generate highly detailed and enhanced images from standard medical scan images, potentially offering doctors to detect anomalies and diseases with greater precision. Furthermore, image generation AI can be used to create alternative versions of medical images to consider what-if scenarios. For example, it can generate virtual images of a patient when they are healthy, which allows for comparing the current actual images with the virtual healthy images, making it possible to provide a diagnosis tailored to the individual patient.
On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions. Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, it may inherently contain biases and errors. Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes. When making critical decisions based on generated images, it is necessary to be able to assess their reliability by properly taking into account the fact that the images were generated by AI. However, to our knowledge, there are no studies that can quantify the reliability of decision-making based on image generation AI.
Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test. When a test image, which may contain an abnormal region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process. By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated. By comparing the input image with the normal image, the anomaly region can be identified. In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomaly regions in the form of $p$-value. The proposed DMAD-test calculates the $p$-values by incorporating the fact that the anomaly region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
In this study, we address this challenge using the statistical hypothesis testing framework. We introduce a statistical inference framework called selective inference (SI), which has gained attention over the past decade in the statistics community as a novel approach for data-driven hypotheses (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014). In SI, statistical inference is performed based on the sampling distribution of the test statistic under the condition that the hypothesis being tested was selected based on the data. Our core idea is to formulate decision-making tasks involving generated images as statistical hypothesis testing problems, and to incorporate SI framework to accurately quantify the reliability of decisions influenced by these generated images.
As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomaly regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1). Initially, a diffusion model is trained exclusively on normal images during the training phase. In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomaly regions. Our proposed statistical test, the Diffusion Model-based Anomaly Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values. Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05) even in finite sample settings.
Related works.
Diffusion models have been effectively utilized in anomaly region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023). In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains abnormal regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomaly regions. Other types of generative AI has also been used for anomaly region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them work on image generation by diffusion models.
Contributions.
Our main contributions in this study are summarized as follows. Our first contribution is the introduction of a statistical testing framework for quantifying reliability in decision-making based on images generated by diffusion models. The second contribution is the implementation of SI for diffusion models, which requires the calculation of the sampling distribution conditional on the diffusion model, necessitating the development of non-trivial computational methodology. The third contribution is to theoretically guarantee the performance of the proposed DMAD-test and demonstrate its performance through numerical experiments and applications in brain imaging diagnostics. The code is available as supplementary material.
2 Diffusion Models
In this section, we briefly explain the diffusion model employed in this study. Given a test image which possibly contain abnormal regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$. Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\mathbf{x}_{0}$, noisy images $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\mathbf{x})$, and approximate the distribution by a parametric model $p_{\theta}(\mathbf{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{(t\geq 1)}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$
where
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\bm{I}).$
By the reproducibility of the Gaussian distribution, $\mathbf{x}_{t}$ can be rewritten by a linear combination of $\mathbf{x}_{0}$ and $\epsilon$, i.e.,
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I}),$ (1)
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}\bm{I})$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon||^{2}_{2}$. Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t},$ (2)
where
$f^{(t)}_{\theta}(\mathbf{x}_{t}):=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}},$ (3)
and
$\sigma_{t}=\eta\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}.$ (4)
Here, $\eta$ is a hyperparameter that controls the randomness in the reverse process. By setting $\eta=1$, we can create new images by stochastic sampling. On the other hand, if we set $\eta=0$, deterministic sampling is used for image generation. By recursively sampling as in (2), we can obtain a reconstructed image of the original input $\mathbf{x}_{0}$.
In practice, the reverse process starts from $\mathbf{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contain individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the abnormal regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023). Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the abnormal region to the normal region. Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost. For the purpose of reducing computational cost, various methods have been proposed. For example, one way is to sample while skipping portions of the sampling trajectory (see Appendix A). The image reconstruction scheme by DDPM is summarized in Algorithm 1.
Algorithm 1 Reconstruction Process
0:  Input image $\mathbf{x}$
1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon_{T^{\prime}}$
2:  for $t=T^{\prime},\dots,1$ do
3:     $f^{(t)}_{\theta}(\mathbf{x}_{t})\leftarrow(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$
4:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t}$
5:  end for
5:  Reconstructed image $\mathbf{x}_{0}$
3 Statistical Test on Generated Images by Diffusion Models
In this section, we formulate the statistical test for detecting anomaly regions using images generated by a trained DDPM model. As shown in Figure 1, anomaly region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomaly regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomaly regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomaly regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomaly regions can be detected by comparing the original test image and its reconstructed one.
Problem Formulation.
We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models. To develop a statistical test, we interpret that an image is considered as a sum of true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and noise component $\bm{\epsilon}\in\mathbb{R}^{n}$. Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint. On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and their covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI. Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
$\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\epsilon},\;\;\;\bm{\epsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (5)
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix, which is known or estimable from external data. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}^{\rm obs}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomaly regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
$E:\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (6)
where absolute value is taken pixel-wise. Anomaly regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomaly region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
$\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid E_{i}(\bm{X})\geq\lambda\right\}.$ (7)
Statistical Inference.
In order to quantify the statistical significance of the anomaly regions detected by a diffusion model, let us consider a statistical test with the following null and alternative hypotheses:
$\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i}~{}~{}~{}\text{v.s.}~{}~{}~{}\mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i},$ (8)
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the anomaly regions and other regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (8) is the difference in mean pixel values between the inside and outside of the anomaly region, i.e.,
$T(\bm{X})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\sum_{i\in\mathcal{M}^{c}_{\bm{X}}}X_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\bm{X},$ (9)
where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}^{c}_{\bm{X}}}$ is the vector that depends on the anomaly region $\mathcal{M}_{\bm{X}}$ and $\mathbf{1}^{n}_{\mathcal{C}}\in\mathcal{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
If we do not account for the fact that the anomaly region is detected by a diffusion model, the distribution of the test statistic would be simply given as $T(\bm{X})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. In this case, the $p$-values defined as
$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\right),$ (10)
would be easily computed by the normality of the test statistic distribution. However, in reality, since the anomaly region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated. Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.
4 Computing Selective $p$-values
In this section, we introduce selective inference (SI) framework for testing images generated by diffusion models and propose a method to perform valid hypothesis test.
4.1 Conditional Distribution of Test Statistics
Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X})$. Then, we consider the sampling distribution of $T(\bm{X})$ conditional on the event that the anomaly region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomaly region $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$, i.e.,
$T(\bm{X})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (11)
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameters that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The nuisance parameter $\mathcal{Q}_{\bm{X}}$ is written as
$\mathcal{Q}_{\bm{X}}=\left(I_{n}-\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\bm{X}.$ (12)
The $p$-value calculated from this conditional sampling distribution is called a selective $p$-value. Specifically, the selective $p$-value is defined as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (13)
where $\mathcal{X}$ is the conditional data space defined as
$\mathcal{X}=\left\{\bm{X}\in\mathbb{R}^{n}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (14)
Due to the conditioning on the nuisance parameter $\mathcal{Q}_{\bm{X}}$, the conditional data space $\mathcal{X}$ can be rewritten as
$\mathcal{X}=\left\{\bm{X}(z)\in\mathbb{R}^{n}\mid\bm{X}(z)=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (15)
where vectors $\bm{a},\bm{b}\in\mathbb{R}^{n}$ are defined as
$\bm{a}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}},\;\bm{b}=\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}},$ (16)
and the region $\mathcal{Z}$ is defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (17)
Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z^{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}+\bm{b}Z$ and $\bm{X}^{\mathrm{obs}}=\bm{a}+\bm{b}z^{\mathrm{obs}}$. Then, the selective $p$-value in (13) is re-written as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (18)
Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution. Once the truncated region $\mathcal{Z}$ is identified, computing the selective $p$-value in (18) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
4.2 Over-Conditioning
To compute the truncated region $\mathcal{Z}$, we employ a divide and conquer approach. It is difficult to directly identify the truncated region $\mathcal{Z}$ due to the complexity of the computational algorithm of the diffusion model. The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022). It is easy to understood that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$. Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomaly region) as the observed one is selected. In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$. By exploiting this, we identify a over-conditioned interval $\mathcal{Z}^{\mathrm{oc}}\subset\mathcal{Z}$.
Identification of $\mathcal{Z}^{\mathrm{oc}}$.
Let us write a polyhedron $\mathcal{P}$ composed of piecewise-linear functions as
$\mathcal{P}_{k}=\left\{\bm{\Delta}_{k}\bm{X}\leq\bm{\delta}_{k}\right\},k\in[K],$ (19)
where $\bm{\Delta}_{k}$ and $\bm{\delta}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions of the $k$-th piecewise-linear function, respectively. Then, a piecewise-linear function $\mathcal{A}(\bm{X})$ is written in the following form:
$\mathcal{A}(\bm{X})=\begin{cases}\bm{\Psi}_{1}\bm{X}+\bm{\psi}_{1}&\text{if }\bm{X}\in\mathcal{P}_{1},\\ \bm{\Psi}_{2}\bm{X}+\bm{\psi}_{2}&\text{if }\bm{X}\in\mathcal{P}_{2},\\ \hskip 28.45274pt\vdots\\ \bm{\Psi}_{K}\bm{X}+\bm{\psi}_{K}&\text{if }\bm{X}\in\mathcal{P}_{K},\end{cases}$ (20)
where $\bm{\Psi}_{k}$ and $\bm{\psi}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions for the $k$-th polyhedron, respectively. Using the notation in (15), since the input image $\bm{X}(z)$ is restricted on a one-dimensional line, each component of the output of $\mathcal{A}$ is written as
$\begin{gathered}\mathcal{A}_{i}(\bm{X}(z))=\begin{cases}\kappa^{\mathcal{A}_{i}}_{1}z+\rho^{\mathcal{A}_{i}}_{1}&\text{if }[L^{\mathcal{A}_{i}}_{1},U^{\mathcal{A}_{i}}_{1}],\\ \kappa^{\mathcal{A}_{i}}_{2}z+\rho^{\mathcal{A}_{i}}_{2}&\text{if }[L^{\mathcal{A}_{i}}_{2},U^{\mathcal{A}_{i}}_{2}],\\ \hskip 28.45274pt\vdots\\ \kappa^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}z+\rho^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}&\text{if }[L^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})},U^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}],\end{cases}\end{gathered}$ (21)
where $K(\mathcal{A}_{i})$ is the number of linear pieces of the piecewise-linear function, and $\kappa^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ and $\rho^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ for $k\in[K(\mathcal{A}_{i})]$ are the coefficient and the constant of the $k$-th polyhedron, respectively. For each $i\in[n]$, there exists $k\in[K(\mathcal{A}_{i})]$ such that $z\in[L^{\mathcal{A}_{i}}_{k},U^{\mathcal{A}_{i}}_{k}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
$\begin{gathered}\coloneqq\begin{cases}\left[\max\left(L^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right),U^{\mathcal{A}_{i}}_{k}\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}>0,\\ \left[L^{\mathcal{A}_{i}}_{k},\min\left(U^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right)\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}<0.\end{cases}\end{gathered}$ (22)
We denote the over-conditioned interval as
$\mathcal{Z}^{\text{oc}}(\bm{a}+\bm{b}z)=\bigcap_{i\in[n]}\left[L^{i}_{z},U^{i}_{z}\right].$ (23)
Piecewise Linearity of Diffusion Models.
We now show that the diffusion model mapping $\mathcal{D}$ and then filtered reconstruction error $E$ can be expressed as a piecewise-linear function of $\bm{X}$. To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below. It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$. To make the reverse process a piecewise-linear function, we employ a U-Net architecture composed of piecewise-linear components such as ReLU activation function and average pooling. Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$. Moreover, since $f^{(t)}_{\theta}(\mathbf{x}_{t})$ in (3) is a composite function of $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$, it is also a piecewise-linear function. By combining them together, we see that $\mathbf{x}_{t-1}$ is written as a piecewise-linear function of $\mathbf{x}_{t}$. Therefore, the entire reconstruction process is a piecewise-linear function since it just repeats the above operation multiple times (see Algorithm 1). As a result, the entire mapping ${\mathcal{D}}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$. Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=E(\bm{X}))$ is piecewise-linear. By exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed.
4.3 Identification of $\mathcal{Z}$ by Parametric Programming
Over-conditioning causes a reduction in power due to excessive conditioning. A technique called Parametric Programming is utilized to explore all intervals along the one-dimensional line, resulting in (17). The truncated region $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{oc}}$ as
$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (24)
The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them. In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z^{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z^{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X})$. An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
Algorithm 2 Selective $p$-value Computation by Parametric Programming
0:  $\mathbf{X^{\mathrm{obs}}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z^{\mathrm{obs}}:=T(\bm{X}^{\mathrm{obs}})$
1:  $\mathcal{Z}\leftarrow\emptyset$
2:  Obtain $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ by (7)
3:  Compute $\bm{a}$, $\bm{b}$ by (17)
4:  $z\leftarrow z_{\mathrm{min}}$
5:  while $z<z_{\mathrm{max}}$ do
6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$ and $\mathcal{M}_{\bm{a}+\bm{b}z}$ by (23) for $z$
7:     if $\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ then
8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$
9:     end if
10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})}+\varepsilon$, where $\varepsilon$ is small positive number.
11:  end while
12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
12:  $p_{\mathrm{selective}}$
5 Experiments
We compared our proposed methods (DMAD-test, DMAD-test-oc) with the other methods: naive method (naive), bonferroni correction (bonferroni), and permutation test (permutation) on type I error rate and power. The details of the methods for comparison are described in Appendix B. We executed the experiment on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
5.1 Numerical Experiments
Experimental Setup.
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. The details of the structure of the diffusion model we used are shown in Appendix C. The diffusion models were trained on data tailored to each experimental settings. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=200$, and the reconstruction was conducted $5$ step samplings. The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In this experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$. We used only normal images for the type I error rate experiments. The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{0},\Sigma)$. we made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only anomaly images. We made 1000 anomaly images $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{\mu},\Sigma)$, where $\mu_{i}=\Delta,\forall i\in\mathcal{S}$, and $\mu_{i}=0,\forall i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomaly region with its position randomly chosen. The image size of the anomaly images was set to 256, with signals $\Delta\in\left\{1,2,3,4\right\}$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $2$. All experiments were conducted under the significance level $\alpha=0.05$. In addition, we conducted robustness experiments against non-Gaussian noise, i.e., $\bm{X}\not\sim\mathcal{N}$. The details are described in Appendix D.
Results.
Figure 2 shows type I error rate. The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$. In contrast, naive and permutation cannot control the type I error rate. Figure 3 shows powers. Since naive and permutation cannot control the type I error rate, their powers are not considered. Among the methods that can control the type I error rate, the proposed method has the highest power. DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
(a) Independence
(b) Correlation
Figure 2: Type I Error Rate comparison
(a) Independence
(b) Correlation
Figure 3: Power comparison
5.2 Real Data Experiments
We used the dataset in (Buda et al., 2019). This dataset includes brain images with tumors as normal images (939 images) and brain images without tumors as anomaly images (941 images). A random selection of 100 images from the normal images was made to estimate the mean and variance. The remaining images were standardized using the estimated mean and variance, and 750 normal images were extracted for training the diffusion model. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=330$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.8$ and the kernel size of the averaging filter to $2$. The results of the DMAD-test and naive are shown in the Figure 4. The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors. The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for anomaly images (True Positive), indicating that the results are desirable. On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for anomaly images but also for normal images (False Positive), indicating the invalidness of the naive test. This result indicates that the DMAD-test detected anomaly region as statistically significant while avoiding misidentification of normal image as anomaly.
$p_{\rm naive}=0.027$, $p_{\rm selective}=0.866$ Refer to caption $p_{\rm naive}=0.028$, $p_{\rm selective}=0.611$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.530$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.784$ Refer to caption $p_{\rm naive}=0.001$, $p_{\rm selective}=0.937$ Refer to caption $p_{\rm naive}=0.009$, $p_{\rm selective}=0.145$
(a) Brain images without tumor
$p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.032$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.036$
(b) Brain images with tumor
Figure 4: Example results of the proposed DMAD-test and the naive method applied to brain images.
6 Conclusions, Limitations and Future Works
In this study, we proposed a statistical test for anomaly regions in medical images detected by a diffusion model. With the proposed DMAD-test, the false detection probability can be controlled with the desired significance level because statistical inference is properly conducted conditional on the fact that the anomaly regions are identified by a diffusion model. We demonstrated the proposed method on a simple anomaly detection task, but its applicability to other tasks remains unclear. Moreover, as the complexity of the diffusion model architecture increases, so does the computational cost. Therefore, to apply the proposed method to practical problems, it is necessary to improve computational efficiency.
Acknowledgement
This work was partially supported by MEXT KAKENHI (20H00601), JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JST AIP Acceleration Research (JPMJCR21U2), NEDO (JPNP18002, JPNP20006) and RIKEN Center for Advanced Intelligence Project.
Appendix A Accelerated Reverse Processes
Methods for accelerating the reverse process have been proposed in DDPM, DDIM (Song et al., 2022). When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$. In this case, equations (2) and (4) can be rewritten as
$\mathbf{x}_{\tau_{i-1}}=\sqrt{\alpha_{\tau_{i-1}}}\left(\frac{\mathbf{x}_{\tau_{i}}-\sqrt{1-\alpha_{\tau_{i}}}\cdot\epsilon^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})}{\sqrt{\alpha_{\tau_{i}}}}\right)+\sqrt{1-\alpha_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\cdot\epsilon_{\theta}^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})+\sigma_{\tau_{i}}\epsilon_{\tau_{i}},$ (25)
where
$\sigma_{\tau_{i}}=\eta\sqrt{(1-\alpha_{\tau_{i-1}})/(1-\alpha_{\tau_{i}})}\sqrt{1-\alpha_{\tau_{i}}/\alpha_{\tau_{i-1}}}.$ (26)
Therefore, piecewise-linearity is preserved, making the proposed method DMAD-test applicable.
Appendix B Comparison Methods for Numerical Experiments
We compared our proposed method with the following methods:
  • DMAD-test: The proposed method uses the parametric programming.
  • DMAD-test-oc: The proposed method uses over-conditioning.
  • naive: The naive method. This method uses a conventional $z$-test without any conditioning. The naive $p$-value is calculated as
    $p_{\mathrm{naive}}=\mathbb{P}_{\mathbb{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\right).$
  • bonferroni: To control the type I error rate, this method applies the bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
    $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
  • permutation: This method uses a permutation test with the steps outlined below:
    • Calculate the observed test statistic $z^{\mathrm{obs}}$ by applying the observated image $\bm{X}^{\mathrm{obs}}$ to the diffusion model.
    • For each $i=1,\ldots,B$, compute the test statistic $z^{(i)}$ by applying the permuted image $\bm{X}^{(i)}$ to the diffusion model, where $B$ represents the total number of permutations, set to 1,000 in our experiments.
    • The permutation $p$-value is then determined as
      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z^{\mathrm{obs}}|\},$
      where $\bm{1}\{\cdot\}$ denotes the indicator function.
This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
Appendix C Structure of the U-Net
Figure 5 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks. The U-Net employs a dense layer for embedding time $t$.
Figure 5: The architecture of the U-Net
Appendix D Robustness Experiments
To evaluate the robustness of our proposed method against non-Gaussian noise, we conducted experiments using various non-Gaussian distributions with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$.
D.1 Experimental Setup
We considered the following non-Gaussian distributions, each with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
  • skew-norm: Skew normal distribution.
  • exponnorm: Exponentially modified normal distribution.
  • gennormsteep: Generalized normal distribution.(with a steepness parameter $\beta$ to be steeper than the normal distribution, i.e., $\beta<2$).
  • gennormflat: Generalized normal distribution (with a fatness parameter $\beta$ to be fatter than the normal distribution, i.e., $\beta>2$).
  • t: Student’s $t$-distribution.
Note that these distributions are standardized in the experiments. Figure 6 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$. We conducted experiments for each distribution family and each 1-Wasserstein distance $1000$ times to calculate the type I error rate. The significance levels $\alpha$ were set to $0.05$ and $0.10$.
Figure 6: Non-Gaussian distributions with $d=0.04$
D.2 Results
The results of the robustness experiments are shown in Figure 7. Our proposed method, DMAD-test, maintains good performance on type I error rate for all the considered distributions.
(a) $\alpha=0.05$
(b) $\alpha=0.10$
Figure 7: Type I error rate for non-Gaussian distributions

### v3

@@
-Statistical Test on Diffusion Model-based Generated Images by Selective Inference
+Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference
 Abstract
-AI technology for generating images, such as diffusion models, has advanced rapidly.
-However, there is no established framework for quantifying the reliability of AI-generated images, which hinders their use in critical decision-making tasks, such as medical image diagnosis.
-In this study, we propose a method to quantify the reliability of decision-making tasks that rely on images produced by diffusion models within a statistical testing framework.
-The core concept of our statistical test involves using a selective inference framework, in which the statistical test is conducted under the condition that the images are produced by a trained diffusion model.
-As a case study, we study a diffusion model-based anomaly detection task for medical images.
-With our approach, the statistical significance of medical image diagnostic outcomes can be quantified in terms of a $p$-value, enabling decision-making with a controlled error rate.
+Advancements in AI image generation, particularly diffusion models, have progressed rapidly.
+However, the absence of an established framework for quantifying the reliability of AI-generated images hinders their use in critical decision-making tasks, such as medical image diagnosis.
+In this study, we address the task of detecting anomalous regions in medical images using diffusion models and propose a statistical method to quantify the reliability of the detected anomalies.
+The core concept of our method involves a selective inference framework, wherein statistical tests are conducted under the condition that the images are produced by a diffusion model.
+With our approach, the statistical significance of anomaly detection results can be quantified in the form of a $p$-value, enabling decision-making with controlled error rates, as is standard in medical practice.
 We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
@@
 On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions.
-Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, it may inherently contain biases and errors.
+Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, they may inherently contain biases and errors.
 Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes.
@@
 Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test.
-When a test image, which may contain an abnormal region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process.
+When a test image, which may contain an anomalous region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process.
 By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated.
-By comparing the input image with the normal image, the anomaly region can be identified.
-In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomaly regions in the form of $p$-value.
-The proposed DMAD-test calculates the $p$-values by incorporating the fact that the anomaly region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
+By comparing the input image with the normal image, the anomalous region can be identified.
+In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomalous regions in the form of $p$-value.
+The DMAD-test calculates the $p$-values by incorporating the fact that the anomalous region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
 In this study, we address this challenge using the statistical hypothesis testing framework.
@@
 Our core idea is to formulate decision-making tasks involving generated images as statistical hypothesis testing problems, and to incorporate SI framework to accurately quantify the reliability of decisions influenced by these generated images.
-As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomaly regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1).
+As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomalous regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1).
 Initially, a diffusion model is trained exclusively on normal images during the training phase.
-In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomaly regions.
-Our proposed statistical test, the Diffusion Model-based Anomaly Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values.
-Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05) even in finite sample settings.
-Related works.
-Diffusion models have been effectively utilized in anomaly region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023).
+In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomalous regions.
+Our proposed statistical test, the Diffusion Model-based Anomalous Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values.
+Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05).
+Related work.
+Diffusion models have been effectively utilized in anomalous region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023).
 In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022).
@@
 In the test phase, the model attempts to reconstruct a new test image.
-If the image contains abnormal regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions.
-The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomaly regions.
-Other types of generative AI has also been used for anomaly region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
+If the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions.
+The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions.
+Other types of generative AI has also been used for anomalous region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
 SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021).
@@
 To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event.
-To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them work on image generation by diffusion models.
+To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them works on image generation by diffusion models.
 Contributions.
@@
 In this section, we briefly explain the diffusion model employed in this study.
-Given a test image which possibly contain abnormal regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image.
+Given a test image which possibly contain anomalous regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image.
 The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
-In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$.
+In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$.
 Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value.
@@
 Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
-$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{(t\geq 1)}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$
-where
-$q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\bm{I}).$
+$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad\text{where}\;\;q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}I).$
 By the reproducibility of the Gaussian distribution, $\mathbf{x}_{t}$ can be rewritten by a linear combination of $\mathbf{x}_{0}$ and $\epsilon$, i.e.,
-$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I}),$ (1)
+$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},I),$ (1)
 where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
-In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}\bm{I})$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$.
+In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$.
 Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$.
-In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon||^{2}_{2}$.
+In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon_{t}||^{2}_{2}$.
 Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
@@
 and
-$\sigma_{t}=\eta\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}.$ (4)
+$\sigma_{t}=\eta\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}.$ (4)
 Here, $\eta$ is a hyperparameter that controls the randomness in the reverse process.
@@
 The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image.
-Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the abnormal regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023).
-Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the abnormal region to the normal region.
+Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023).
+Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the anomalous region to the normal region.
 Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost.
@@
 0:  Input image $\mathbf{x}$
-1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon_{T^{\prime}}$
+1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
 2:  for $t=T^{\prime},\dots,1$ do
@@
 3 Statistical Test on Generated Images by Diffusion Models
-In this section, we formulate the statistical test for detecting anomaly regions using images generated by a trained DDPM model.
-As shown in Figure 1, anomaly region detection by diffusion models is conducted as follows.
+In this section, we formulate the statistical test for detecting anomalous regions using images generated by a trained DDPM model.
+As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows.
 First, in the training phase, the diffusion model is trained only on normal images.
-Then, in the test phase, we feed a test image which might contain anomaly regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$.
+Then, in the test phase, we feed a test image which might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$.
 By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image.
-If the image does not contain anomaly regions, the reconstructed image is expected to be similar to the original test image.
-On the other hand, if the image contains anomaly regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions.
-Therefore, the anomaly regions can be detected by comparing the original test image and its reconstructed one.
+If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image.
+On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions.
+Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
 Problem formulation.
 We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models.
-To develop a statistical test, we interpret that an image is considered as a sum of true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and noise component $\bm{\epsilon}\in\mathbb{R}^{n}$.
+To develop a statistical test, we interpret an image as a sum of a true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and a noise component $\bm{\varepsilon}\in\mathbb{R}^{n}$.
+We emphasize that the noise component $\bm{\varepsilon}$ should not be confused with the noise $\epsilon$ used in the forward process.
 Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint.
@@
 Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
-$\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\epsilon},\;\;\;\bm{\epsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (5)
-where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix, which is known or estimable from external data.
-In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}^{\rm obs}$.
+$\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\varepsilon},\;\;\;\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (5)
+where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix.
+In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}_{\rm obs}$.
 Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$.
 The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error.
-When identifying anomaly regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise.
+When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise.
 In this study, we simply used an averaging filter.
@@
 where absolute value is taken pixel-wise.
-Anomaly regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$.
-Specifically, we define the anomaly region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
+Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$.
+Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
 $\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid E_{i}(\bm{X})\geq\lambda\right\}.$ (7)
 Statistical inference.
-In order to quantify the statistical significance of the anomaly regions detected by a diffusion model, let us consider a statistical test with the following null and alternative hypotheses:
-$\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i}~{}~{}~{}\text{v.s.}~{}~{}~{}\mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i},$ (8)
-where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the anomaly regions and other regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different.
-A reasonable test statistic for the statistical test in (8) is the difference in mean pixel values between the inside and outside of the anomaly region, i.e.,
-$T(\bm{X})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\sum_{i\in\mathcal{M}^{c}_{\bm{X}}}X_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\bm{X},$ (9)
-where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}^{c}_{\bm{X}}}$ is the vector that depends on the anomaly region $\mathcal{M}_{\bm{X}}$ and $\mathbf{1}^{n}_{\mathcal{C}}\in\mathcal{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
-If we do not account for the fact that the anomaly region is detected by a diffusion model, the distribution of the test statistic would be simply given as $T(\bm{X})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$.
+In order to quantify the statistical significance of the anomalous regions detected by a diffusion model, we consider the concrete example of two-sample test.
+Let us define an $n$-dimensional reference input vector,
+$\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),$ (8)
+where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image and the $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component.
+Then, we consider the following null and alternative hypotheses:
+$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i}\\ \text{v.s.}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i},\end{gathered}$ (9)
+where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the anomalous regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different.
+A reasonable test statistic for the statistical test in (9) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{X}}$, i.e.,
+$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X^{\mathrm{ref}}_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (10)
+where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}\in\mathbb{R}^{2n}$ is the vector that depends on the anomalous region $\mathcal{M}_{\bm{X}}$, defined as
+$\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (11)
+where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
+If we do not account for the fact that the anomalous region is detected by a diffusion model, the distribution of the test statistic would be simply given as
+$T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}),\;\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.$ (12)
 In this case, the $p$-values defined as
-$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\right),$ (10)
+$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\right),$ (13)
 would be easily computed by the normality of the test statistic distribution.
-However, in reality, since the anomaly region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated.
+However, in reality, since the anomalous region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated.
 Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.
@@
 4.1 Conditional Distribution of Test Statistics
-Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X})$.
-Then, we consider the sampling distribution of $T(\bm{X})$ conditional on the event that the anomaly region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomaly region $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$, i.e.,
-$T(\bm{X})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (11)
-In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameters that is independent of the test statistic.
+Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$.
+Then, we consider the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$ conditional on the event that the anomalous region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomalous region $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$, i.e.,
+$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (14)
+In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic.
 As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space.
-The nuisance parameter $\mathcal{Q}_{\bm{X}}$ is written as
-$\mathcal{Q}_{\bm{X}}=\left(I_{n}-\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\bm{X}.$ (12)
+The nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
+$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}.$ (15)
 The $p$-value calculated from this conditional sampling distribution is called a selective $p$-value. Specifically, the selective $p$-value is defined as
-$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (13)
+$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (16)
 where $\mathcal{X}$ is the conditional data space defined as
-$\mathcal{X}=\left\{\bm{X}\in\mathbb{R}^{n}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (14)
+[…]
 Due to the conditioning on the nuisance parameter $\mathcal{Q}_{\bm{X}}$, the conditional data space $\mathcal{X}$ can be rewritten as
-$\mathcal{X}=\left\{\bm{X}(z)\in\mathbb{R}^{n}\mid\bm{X}(z)=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (15)
-where vectors $\bm{a},\bm{b}\in\mathbb{R}^{n}$ are defined as
-$\bm{a}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}},\;\bm{b}=\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}},$ (16)
+$\mathcal{X}=\left\{\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}\in\mathbb{R}^{2n}\middle|\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (18)
+where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$, and $\bm{c}_{1:n}$ represents a vector composed of the first $n$ elements of the vector $\bm{c}$.
+The vectors $\bm{a},\bm{b}\in\mathbb{R}^{2n}$ are defined as
+$\bm{a}=\mathcal{Q}_{\bm{X}_{\mathrm{obs}}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}},$ (19)
 and the region $\mathcal{Z}$ is defined as
-$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (17)
-Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z^{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}+\bm{b}Z$ and $\bm{X}^{\mathrm{obs}}=\bm{a}+\bm{b}z^{\mathrm{obs}}$.
+$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (20)
+Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z_{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}_{1:n}+\bm{b}_{1:n}Z$ and $\bm{X}_{\mathrm{obs}}=\bm{a}_{1:n}+\bm{b}_{1:n}z_{\mathrm{obs}}$.
 Then, the selective $p$-value in (16) is re-written as
-$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (18)
-Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$.
+$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (21)
+Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}})$.
 Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution.
@@
 The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022).
-It is easy to understood that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$.
-Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomaly region) as the observed one is selected.
+It is easy to understand that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$.
+Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomalous region) as the observed one is selected.
 In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$.
@@
 The truncated region $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{oc}}$ as
-$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (24)
+$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (27)
 The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them.
-In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z^{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z^{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X})$.
+In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z_{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z_{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$.
 An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
 Algorithm 2 Selective $p$-value Computation by Parametric Programming
-0:  $\mathbf{X^{\mathrm{obs}}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z^{\mathrm{obs}}:=T(\bm{X}^{\mathrm{obs}})$
+0:  $\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z_{\mathrm{obs}}:=T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})$
 1:  $\mathcal{Z}\leftarrow\emptyset$
-2:  Obtain $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ by (7)
+2:  Obtain $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ by (7)
 3:  Compute $\bm{a}$, $\bm{b}$ by (20)
@@
 5:  while $z<z_{\mathrm{max}}$ do
-6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$ and $\mathcal{M}_{\bm{a}+\bm{b}z}$ by (23) for $z$
-7:     if $\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ then
-8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$
+6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (26) for $z$
+7:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ then
+8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$
 9:     end if
-10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})}+\varepsilon$, where $\varepsilon$ is small positive number.
+10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)}+\gamma$, where $\gamma$ is small positive number.
 11:  end while
-12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
+12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
 12:  $p_{\mathrm{selective}}$
@@
 Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$.
-The details of the structure of the diffusion model we used are shown in Appendix C.
-The diffusion models were trained on data tailored to each experimental settings.
-The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=200$, and the reconstruction was conducted $5$ step samplings.
+In the type I error rate experiments, we used only normal images.
+The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$.
+We made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$.
+In the power experiments, we used only abnormal images.
+We generated 1000 abnormal images $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{\mu},\Sigma)$.
+The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen.
+The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$.
+In all experiments, we made the synthetic dataset for 1000 reference images to follow $\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$.
+The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$.
+All experiments were conducted under the significance level $\alpha=0.05$.
+The details of the architecture of the diffusion model we used are shown in Appendix C.
+The diffusion models were trained on the normal images from the synthetic dataset.
+The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings.
 The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear.
-In this experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$.
-We used only normal images for the type I error rate experiments.
-The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{0},\Sigma)$.
-we made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$.
-In the power experiments, we used only anomaly images.
-We made 1000 anomaly images $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{\mu},\Sigma)$, where $\mu_{i}=\Delta,\forall i\in\mathcal{S}$, and $\mu_{i}=0,\forall i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomaly region with its position randomly chosen.
-The image size of the anomaly images was set to 256, with signals $\Delta\in\left\{1,2,3,4\right\}$.
-The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $2$.
-All experiments were conducted under the significance level $\alpha=0.05$.
-In addition, we conducted robustness experiments against non-Gaussian noise, i.e., $\bm{X}\not\sim\mathcal{N}$.
-The details are described in Appendix D.
+In all experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$.
+In addition, we conducted robustness experiments against non-Gaussian noise.
+To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$.
+We considered the following non-Gaussian distributions with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
+  • Skew normal distribution family (SND).
+  • Exponentially modified gaussian distribution family (EMG).
+  • Generalized normal distribution family (GND) with a shape parameter $\beta$. This distribution family can be steeper than the normal distribution (i.e., $\beta<2$).
+  • Student’s $t$-distribution family ($t$-distribution).
+Note that these distributions are standardized in the experiments.
+Figure 2 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$.
+We run $1000$ trials for each distribution family and each 1-Wasserstein distance to calculate the type I error rate.
+The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$.
+Figure 2: Non-Gaussian distributions with $d=0.04$
 Results.
-Figure 2 shows type I error rate.
+Figure 4 shows type I error rate comparison.
 The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$.
 In contrast, naive and permutation cannot control the type I error rate.
-Figure 3 shows powers.
+Figure 4 shows power comparison.
 Since naive and permutation cannot control the type I error rate, their powers are not considered.
@@
 DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
+Figure 5 shows the results of the robustness experiments.
+DMAD-test maintains good performance on the type I error rate for all the considered distribution families.
 (a) Independence
@@
 Figure 4: Power comparison
+(a) SND
+(b) EMG
+(c) GND
+(d) $t$-distribution
+Figure 5: Type I Error Rate for Non-Gaussian distribution families
 5.2 Real Data Experiments
-We used the dataset in (Buda et al., 2019).
-This dataset includes brain images with tumors as normal images (939 images) and brain images without tumors as anomaly images (941 images).
+We examined the medical image experiment using the dataset extracted from the dataset in Buda et al.
+(2019).
+This dataset includes brain images without tumors as normal images (941 images) and brain images with tumors as abnormal images (939 images).
 A random selection of 100 images from the normal images was made to estimate the mean and variance.
-The remaining images were standardized using the estimated mean and variance, and 750 normal images were extracted for training the diffusion model.
-The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=330$, with reconstruction performed through 5 step samplings.
-We set the threshold $\lambda=0.8$ and the kernel size of the averaging filter to $2$.
+The remaining images were standardized using the estimated mean and variance, and 329 normal images were extracted for training the diffusion model.
+The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings.
+We set the threshold $\lambda=0.9$ and the kernel size of the averaging filter to $3$.
 The results of the DMAD-test and naive are shown in the Figure 6.
 The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors.
-The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for anomaly images (True Positive), indicating that the results are desirable.
-On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for anomaly images but also for normal images (False Positive), indicating the invalidness of the naive test.
-This result indicates that the DMAD-test detected anomaly region as statistically significant while avoiding misidentification of normal image as anomaly.
-$p_{\rm naive}=0.027$, $p_{\rm selective}=0.866$ Refer to caption $p_{\rm naive}=0.028$, $p_{\rm selective}=0.611$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.530$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.784$ Refer to caption $p_{\rm naive}=0.001$, $p_{\rm selective}=0.937$ Refer to caption $p_{\rm naive}=0.009$, $p_{\rm selective}=0.145$
-(a) Brain images without tumor
-$p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.032$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.036$
-(b) Brain images with tumor
-Figure 4: Example results of the proposed DMAD-test and the naive method applied to brain images.
-6 Conclusions, Limitations and Future Works
-In this study, we proposed a statistical test for anomaly regions in medical images detected by a diffusion model.
-With the proposed DMAD-test, the false detection probability can be controlled with the desired significance level because statistical inference is properly conducted conditional on the fact that the anomaly regions are identified by a diffusion model.
-We demonstrated the proposed method on a simple anomaly detection task, but its applicability to other tasks remains unclear.
-Moreover, as the complexity of the diffusion model architecture increases, so does the computational cost.
-Therefore, to apply the proposed method to practical problems, it is necessary to improve computational efficiency.
+The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for abnormal images (True Positive), indicating that the results are desirable.
+On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for abnormal images but also for normal images (False Positive), indicating the invalidness of the naive test.
+This result indicates that the DMAD-test detected anomalous regions as statistically significant while avoiding misidentification of normal image as anomaly.
+$p_{\rm naive}=0.019$, $p_{\rm selective}=0.643$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.379$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.290$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.822$ Refer to caption $p_{\rm naive}=0.018$, $p_{\rm selective}=0.373$ Refer to caption $p_{\rm naive}=0.024$, $p_{\rm selective}=0.932$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.225$
+(a) Brain images without tumors
+$p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.002$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.021$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.028$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.027$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.016$
+(b) Brain images with tumors
+Figure 6: An example of the results for applying the proposed DMAD-test and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to brain images.
+6 Conclusions
+In this study, we proposed a statistical test for anomalous regions in medical images detected by a diffusion model.
+With the proposed DMAD-test, the false detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by a diffusion model.
+We believe this study marks a step toward bridging the gap between generative AI and rigorous statistical inference in medical imaging analysis.
 Acknowledgement
@@
   • naive: The naive method. This method uses a conventional $z$-test without any conditioning. The naive $p$-value is calculated as
-    $p_{\mathrm{naive}}=\mathbb{P}_{\mathbb{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\right).$
+    $p_{\mathrm{naive}}=\mathbb{P}_{\mathbb{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\right).$
   • bonferroni: To control the type I error rate, this method applies the bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
@@
   • permutation: This method uses a permutation test with the steps outlined below:
-    • Calculate the observed test statistic $z^{\mathrm{obs}}$ by applying the observated image $\bm{X}^{\mathrm{obs}}$ to the diffusion model.
+    • Calculate the observed test statistic $z_{\mathrm{obs}}$ by applying the observated image $\bm{X}_{\mathrm{obs}}$ to the diffusion model.
     • For each $i=1,\ldots,B$, compute the test statistic $z^{(i)}$ by applying the permuted image $\bm{X}^{(i)}$ to the diffusion model, where $B$ represents the total number of permutations, set to 1,000 in our experiments.
     • The permutation $p$-value is then determined as
-      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z^{\mathrm{obs}}|\},$
+      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z_{\mathrm{obs}}|\},$
       where $\bm{1}\{\cdot\}$ denotes the indicator function.
 This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
-Appendix C Structure of the U-Net
-Figure 5 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks. The U-Net employs a dense layer for embedding time $t$.
+Appendix C Architecture of the U-Net
+Figure 7 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks.
 Figure 7: The architecture of the U-Net
-Appendix D Robustness Experiments
-To evaluate the robustness of our proposed method against non-Gaussian noise, we conducted experiments using various non-Gaussian distributions with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$.
-D.1 Experimental Setup
-We considered the following non-Gaussian distributions, each with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
-  • skew-norm: Skew normal distribution.
-  • exponnorm: Exponentially modified normal distribution.
-  • gennormsteep: Generalized normal distribution.(with a steepness parameter $\beta$ to be steeper than the normal distribution, i.e., $\beta<2$).
-  • gennormflat: Generalized normal distribution (with a fatness parameter $\beta$ to be fatter than the normal distribution, i.e., $\beta>2$).
-  • t: Student’s $t$-distribution.
-Note that these distributions are standardized in the experiments.
-Figure 6 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$.
-We conducted experiments for each distribution family and each 1-Wasserstein distance $1000$ times to calculate the type I error rate.
-The significance levels $\alpha$ were set to $0.05$ and $0.10$.
-Figure 6: Non-Gaussian distributions with $d=0.04$
-D.2 Results
-The results of the robustness experiments are shown in Figure 7. Our proposed method, DMAD-test, maintains good performance on type I error rate for all the considered distributions.
-(a) $\alpha=0.05$
-(b) $\alpha=0.10$
-Figure 7: Type I error rate for non-Gaussian distributions

### v4

@@
-Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference
+Quantifying Statistical Significance in Diffusion-Based Anomaly Localization via Selective Inference
 Abstract
-Advancements in AI image generation, particularly diffusion models, have progressed rapidly.
-However, the absence of an established framework for quantifying the reliability of AI-generated images hinders their use in critical decision-making tasks, such as medical image diagnosis.
-In this study, we address the task of detecting anomalous regions in medical images using diffusion models and propose a statistical method to quantify the reliability of the detected anomalies.
-The core concept of our method involves a selective inference framework, wherein statistical tests are conducted under the condition that the images are produced by a diffusion model.
-With our approach, the statistical significance of anomaly detection results can be quantified in the form of a $p$-value, enabling decision-making with controlled error rates, as is standard in medical practice.
-We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
+Anomaly localization in images—identifying regions that deviate from expected patterns—is vital in applications such as medical diagnosis and industrial inspection.
+A recent trend is the use of image generation models in anomaly localization, where these models generate normal-looking counterparts of anomalous images, thereby allowing flexible and adaptive anomaly localization.
+However, these methods inherit the uncertainty and bias implicitly embedded in the employed generative model, raising concerns about the reliability.
+To address this, we propose a statistical framework based on selective inference to quantify the significance of detected anomalous regions.
+Our method provides $p$-values to assess the false positive detection rates, providing a principled measure of reliability.
+As a proof of concept, we consider anomaly localization using a diffusion model and its applications to medical diagnoses and industrial inspections.
+The results indicate that the proposed method effectively controls the risk of false positive detection, supporting its use in high-stakes decision-making tasks.
 1 Introduction
-Advances in image generation AI, such as diffusion models, have been remarkable (Song and Ermon, 2019).
-They can generate highly realistic and detailed images, which leads to innovations in various tasks across various fields.
-For example, image generation AI significantly enhances medical image diagnosis by improving accuracy and efficiency.
-It can generate highly detailed and enhanced images from standard medical scan images, potentially offering doctors to detect anomalies and diseases with greater precision.
-Furthermore, image generation AI can be used to create alternative versions of medical images to consider what-if scenarios.
-For example, it can generate virtual images of a patient when they are healthy, which allows for comparing the current actual images with the virtual healthy images, making it possible to provide a diagnosis tailored to the individual patient.
-On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions.
-Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, they may inherently contain biases and errors.
-Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes.
-When making critical decisions based on generated images, it is necessary to be able to assess their reliability by properly taking into account the fact that the images were generated by AI.
-However, to our knowledge, there are no studies that can quantify the reliability of decision-making based on image generation AI.
-Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test.
-When a test image, which may contain an anomalous region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process.
-By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated.
-By comparing the input image with the normal image, the anomalous region can be identified.
-In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomalous regions in the form of $p$-value.
-The DMAD-test calculates the $p$-values by incorporating the fact that the anomalous region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
-In this study, we address this challenge using the statistical hypothesis testing framework.
-We introduce a statistical inference framework called selective inference (SI), which has gained attention over the past decade in the statistics community as a novel approach for data-driven hypotheses (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014).
-In SI, statistical inference is performed based on the sampling distribution of the test statistic under the condition that the hypothesis being tested was selected based on the data.
-Our core idea is to formulate decision-making tasks involving generated images as statistical hypothesis testing problems, and to incorporate SI framework to accurately quantify the reliability of decisions influenced by these generated images.
-As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomalous regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1).
-Initially, a diffusion model is trained exclusively on normal images during the training phase.
-In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomalous regions.
-Our proposed statistical test, the Diffusion Model-based Anomalous Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values.
-Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05).
-Related work.
-Diffusion models have been effectively utilized in anomalous region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023).
-In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022).
+Anomaly localization using image generation models, particularly diffusion models, has shown great promise across diverse domains such as medical diagnosis and industrial inspection Li et al.
+(2023); Iqbal et al.
+(2023); Lu et al.
+(2023); Zhang et al.
+(2023); Fontanella et al.
+(2024); Tebbe and Tayyub (2024); Sheng et al.
+(2024).
+These models reconstruct a normal-looking version of an input image, and differences between the input and the reconstruction highlight potential anomalies.
+Compared to traditional methods, generative approaches are highly suitable for settings where annotations for anomalous regions are unavailable.
+Moreover, generative approaches can flexibly handle heterogeneity by adapting to individual images–e.g., patient-specific characteristics in medical diagnosis and product-specific traits in industrial inspection.
+Among various image generation models, diffusion models, in particular, offer high fidelity and stability, outperforming other methods in image quality and anomaly localization.
+While generative approaches offer powerful and flexible capabilities for anomaly localization, a major concern is that the inherent uncertainty and bias in generative models can affect localization performance Fithian et al.
+(2014); Taylor and Tibshirani (2015); Lee et al.
+(2016); Duy and Takeuchi (2022); Miwa et al.
+(2023); Shiraishi et al.
+(2024).
+These models are trained on specific datasets composed of normal images, and the quality of the generated normal-looking images depends heavily on the distribution of the dataset and how well the model has learned the underlying distribution.
+As a result, uncertainties or biases in the dataset or training process can cause incorrect reconstructions, leading to inaccurate localizations and misidentification of anomalies.
+Such risks are especially critical in high-stakes domains such as medical diagnosis and industrial inspection, where even minor errors can have serious consequences.
+Therefore, it is essential to incorporate rigorous uncertainty quantification framework and statistical safeguards to ensure reliable deployment in critical applications.
+To address this issue, we propose a statistical testing framework based on Selective Inference (SI) to assess the statistical significance of detected anomalies.
+SI has recently emerged as a promising approach for conducting statistical inference on hypotheses that are selected based on observed data.
+In this framework, inference is performed using the conditional distribution given the selection event, thereby accounting for the uncertainty and bias associated with the hypothesis selection.
+Following this SI framework, our key idea is to apply statistical testing to detected anomalies, conditioned on the fact that the anomalous regions were identified using a specific generative model.
+This allows us to quantify the statistical significance of detected anomalies as a valid $p$-value, providing a rigorous estimate of the false positive rate and offering a principled metric for reliability.
+In this paper, as a proof of concept for the proposed statistical testing framework, we focus on a standard denoising diffusion probabilistic model (DDPM) Ho et al.
+(2020); Song et al.
+(2022) among various diffusion-based anomaly detection (AD) methods , and its applications to medical diagnostics and industrial inspections.
+However, the proposed framework is readily generalizable to a broader range of diffusion model architectures and is applicable to semi-supervised AD problems in other domains.
+Related works
+Diffusion models have been effectively utilized in anomaly localization problems Pinaya et al.
+(2022); Fontanella et al.
+(2024); Wyatt et al.
+(2022); Mousakhan et al.
+(2023).
+In this context, the DDPM is commonly used Ho et al.
+(2020); Song et al.
+(2022).
 During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise.
@@
 The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions.
-Other types of generative AI has also been used for anomalous region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
-SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021).
-Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022).
+Other types of generative AI has also been used for anomalous region detection task Baur et al.
+(2021); Chen and Konukoglu (2018); Chow et al.
+(2020); Jana et al.
+(2022).
+SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm Lee and Taylor (2014); Lee et al.
+(2016); Tibshirani et al.
+(2016), and then extended to more complex feature selection methods Yang et al.
+(2016); Suzumura et al.
+(2017); Hyun et al.
+(2018); Rügamer and Greven (2020); Das et al.
+(2021).
+Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks Chen and Bien (2020); Tsukurimichi et al.
+(2021); Tanizaki et al.
+(2020); Duy et al.
+(2022); Le Duy et al.
+(2024); Lee et al.
+(2015); Gao et al.
+(2022); Duy et al.
+(2020); Jewell et al.
+(2022).
 The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data.
 To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event.
-To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them works on image generation by diffusion models.
-Contributions.
-Our main contributions in this study are summarized as follows.
-Our first contribution is the introduction of a statistical testing framework for quantifying reliability in decision-making based on images generated by diffusion models.
-The second contribution is the implementation of SI for diffusion models, which requires the calculation of the sampling distribution conditional on the diffusion model, necessitating the development of non-trivial computational methodology.
-The third contribution is to theoretically guarantee the performance of the proposed DMAD-test and demonstrate its performance through numerical experiments and applications in brain imaging diagnostics.
-The code is available as supplementary material.
-2 Diffusion Models
-In this section, we briefly explain the diffusion model employed in this study.
-Given a test image which possibly contain anomalous regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image.
+To the best of our knowledge, SI was applied to statistical inferences on several deep learning models Duy et al.
+(2022); Miwa et al.
+(2023); Shiraishi et al.
+(2024); Miwa et al.
+(2024), but none of them works on image generation by diffusion models.
+Our contributions
+The main contributions of our study are summarized as follows.
+[^1] We note that our contribution is not the development of a new diffusion-based anomaly localization algorithm, but rather the introduction of a rigorous statistical testing framework designed to quantify the statistical reliability of anomalous regions identified by diffusion-based AD methods.
+  • We propose a novel statistical testing framework to assess the significance of anomaly localization results derived from diffusion model-based methods, offering a principled basis for evaluating the reliability of detected anomalies.
+  • We implement the SI framework for diffusion models by deriving the sampling distribution conditional on the selection event induced by the diffusion model, which requires developing non-trivial computational techniques tailored to the generative sampling process.
+  • We provide theoretical justification for the proposed method and validate its effectiveness through extensive numerical experiments in medical diagnosis and industrial inspection scenarios.
+The results highlight the robustness and practical utility of our method.
+The implementation code for reproducing all experimental results is provided at https://github.com/tkatsuoka/DAL-test
+Figure 1: Schematic illustration of anomaly localization on a brain MRI image dataset using a diffusion model and the proposed DAL-Test.
+When a test image—potentially containing an anomalous region—is fed into a trained diffusion model, its normal-looking version is generated through the forward and reverse processes.
+By comparing the input image with the generated normal-looking version, the anomalous region can be identified.
+We propose the Diffusion-based Anomaly Localization (DAL) Test, which leverages the selective inference framework to compute valid $p$-values that quantify the statistical significance of anomalous regions detected by a diffusion model, based on a test statistic defined over the input and reference images.
+2 Anomaly localization by diffusion models
+This section describes the anomaly localization task based on a diffusion model, which is explored as a proof of concept in this study.
+The process of anomaly localization using generative models can generally be divided into two phases.
+First, during the training phase, a denoising diffusion probabilistic model (DDPM) is trained using a dataset composed exclusively of normal images.
+The model learns the distribution of normal images through two key processes: the diffusion process, in which noise is gradually added to an image, and the reverse diffusion process, in which the original image is reconstructed from noise.
+Through this procedure, the model enhances its capacity to reconstruct normal image structures by acquiring denoising capabilities at each step.
+Next, during the testing phase, the reverse diffusion process is conditionally applied to an unseen input test image.
+In this step, the model reconstructs an image that closely resembles the input but conforms to its learned notion of “normality”, causing anomalous regions to be poorly reproduced.
+An anomaly score is then computed based on the difference between the reconstructed image and the input test image, and the spatial distribution of this score is analyzed to localize anomalies.
+By applying thresholding to the score map, anomalous regions can be clearly identified.
+In this study, for the purpose of proof of concept, we adopt standard denoising diffusion models as our choice of diffusion model Ho et al.
+(2020); Song et al.
+(2022).
+The following outlines the image reconstruction process of a trained DDPM.
+Given a test image which possibly contain anomalous regions, a denoising diffusion model is used to generate the corresponding normal image.
 The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
 In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$.
-Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value.
-Given an original test image $\mathbf{x}_{0}$, noisy images $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps.
-We consider the distribution of the original and noisy test images, which is denoted by $q(\mathbf{x})$, and approximate the distribution by a parametric model $p_{\theta}(\mathbf{x})$ with $\theta$ being the parameters.
+Let $\bm{x}$ be an image represented as a vector with each element corresponding to a pixel value.
+Given an original test image $\bm{x}_{0}$, noisy images $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps.
+We consider the distribution of the original and noisy test images, which is denoted by $q(\bm{x})$, and approximate the distribution by a parametric model $p_{\theta}(\bm{x})$ with $\theta$ being the parameters.
 Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
-$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad\text{where}\;\;q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}I).$
-By the reproducibility of the Gaussian distribution, $\mathbf{x}_{t}$ can be rewritten by a linear combination of $\mathbf{x}_{0}$ and $\epsilon$, i.e.,
-$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},I),$ (1)
+$q(\bm{x}_{1:T}|\bm{x}_{0}):=\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1}),\;\;\text{where}\;\;q(\bm{x}_{t}|\bm{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}I).$ (1)
+By the reproducibility of the Gaussian distribution, $\bm{x}_{t}$ can be rewritten by a linear combination of $\bm{x}_{0}$ and $\epsilon_{t}$, i.e.,
+$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t},\;\;\text{with}\;\;\epsilon_{t}\sim\mathcal{N}(\mathbf{0},I),$ (2)
 where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
-In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$.
-Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$.
-In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon_{t}||^{2}_{2}$.
-Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
-$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t},$ (2)
+In the reverse process, a parametric model in the form of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\mu_{\theta}(\bm{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\bm{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$.
+Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$.
+In DDPM Ho et al.
+(2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\bm{x}_{t})-\epsilon_{t}||^{2}_{2}$.
+Based on (2), given a noisy image $\bm{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\bm{x}_{t-1}$ is obtained as
+$\bm{x}_{t-1}=\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\bm{x}_{t})+\sigma_{t}\epsilon_{t},$ (3)
 where
-$f^{(t)}_{\theta}(\mathbf{x}_{t}):=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}},$ (3)
+$f^{(t)}_{\theta}(\bm{x}_{t}):=(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}},$ (4)
 and
@@
 On the other hand, if we set $\eta=0$, deterministic sampling is used for image generation.
-By recursively sampling as in (2), we can obtain a reconstructed image of the original input $\mathbf{x}_{0}$.
-In practice, the reverse process starts from $\mathbf{x}_{T^{\prime}}$ with $T^{\prime}<T$.
-Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contain individual information of the original input image.
+By recursively sampling as in (3), we can obtain a reconstructed image of the original input $\bm{x}_{0}$.
+In practice, the reverse process starts from $\bm{x}_{T^{\prime}}$ with $T^{\prime}<T$.
+Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contains individual information of the original input image.
 The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image.
-Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023).
-Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the anomalous region to the normal region.
-Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost.
-For the purpose of reducing computational cost, various methods have been proposed.
-For example, one way is to sample while skipping portions of the sampling trajectory (see Appendix A).
-The image reconstruction scheme by DDPM is summarized in Algorithm 1.
+Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions Ho et al.
+(2020); Mousakhan et al.
+(2023).
 Algorithm 1 Reconstruction Process
-0:  Input image $\mathbf{x}$
-1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
+0:  Input image $\bm{x}$
+1:  $\bm{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\bm{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
 2:  for $t=T^{\prime},\dots,1$ do
-3:     $f^{(t)}_{\theta}(\mathbf{x}_{t})\leftarrow(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$
-4:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t}$
+3:     $f^{(t)}_{\theta}(\bm{x}_{t})\leftarrow(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}}$
+4:     $\bm{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t})+\sigma_{t}\epsilon_{t}$
 5:  end for
-5:  Reconstructed image $\mathbf{x}_{0}$
-3 Statistical Test on Generated Images by Diffusion Models
-In this section, we formulate the statistical test for detecting anomalous regions using images generated by a trained DDPM model.
+5:  Reconstructed image $\bm{x}_{0}$
+3 Statistical test for diffusion-based anomalous localization
+In this section, we formulate the statistical test for detecting anomalous regions by a trained diffusion model.
 As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows.
@@
 Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
-Problem formulation.
+Problem formulation
 We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models.
@@
 where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix.
-In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}_{\rm obs}$.
-Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$.
+In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image vector is denoted as $\bm{x}$.
+Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}\colon\mathbb{R}^{n}\ni\bm{X}\to\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$.
 The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error.
@@
 In this study, we simply used an averaging filter.
-Let us denote the averaging filter as $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
+Let us denote the averaging filter as $\mathcal{F}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$.
 Then, the process of obtaining the (filtered) reconstruction error is written as
-$E:\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (6)
+$\mathcal{E}\colon\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (7)
 where absolute value is taken pixel-wise.
-Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$.
-Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
-$\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid E_{i}(\bm{X})\geq\lambda\right\}.$ (7)
-Statistical inference.
-In order to quantify the statistical significance of the anomalous regions detected by a diffusion model, we consider the concrete example of two-sample test.
+Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $\mathcal{E}_{i}(\bm{X})$ for each pixel $i\in[n]$.
+Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in\mathbb{R}^{+}$, i.e.,
+$\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid\mathcal{E}_{i}(\bm{X})\geq\lambda\right\}.$ (8)
+Statistical test for diffusion models
+In order to quantify the statistical significance of the anomalous regions detected by using the diffusion model, we consider the concrete example of two-sample test.
+Note that our method can be extended to other statistical tests using various statistics.
+In the two-sample test, we compare the test image and the randomly chosen reference image in the anomalous region.
 Let us define an $n$-dimensional reference input vector,
-$\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),$ (8)
+$\begin{gathered}\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\text{with}\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),\end{gathered}$ (9)
 where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image and the $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component.
 Then, we consider the following null and alternative hypotheses:
-$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i}\\ \text{v.s.}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i},\end{gathered}$ (9)
-where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the anomalous regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different.
+$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i}\\ ~{}~{}~{}\text{v.s.}~{}~{}~{}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i},\end{gathered}$ (10)
+where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the observed anomalous region $\mathcal{M}_{\bm{x}}$, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different.
 A reasonable test statistic for the statistical test in (10) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
-$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X^{\mathrm{ref}}_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (10)
-where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}\in\mathbb{R}^{2n}$ is the vector that depends on the anomalous region $\mathcal{M}_{\bm{X}}$, defined as
-$\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (11)
+$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}X_{i}-\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}X^{\mathrm{ref}}_{i}=\bm{\nu}^{\top}\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (11)
+where $\bm{\nu}\in\mathbb{R}^{2n}$ denotes a vector that depends on the anomalous region $\mathcal{M}_{\bm{x}}$, and hence on $\bm{x}$ itself, and is defined as
+$\bm{\nu}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{x}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{x}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (12)
 where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
-If we do not account for the fact that the anomalous region is detected by a diffusion model, the distribution of the test statistic would be simply given as
-$T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}),\;\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.$ (12)
-In this case, the $p$-values defined as
-$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\right),$ (13)
-would be easily computed by the normality of the test statistic distribution.
-However, in reality, since the anomalous region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated.
-Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.
-4 Computing Selective $p$-values
-In this section, we introduce selective inference (SI) framework for testing images generated by diffusion models and propose a method to perform valid hypothesis test.
-4.1 Conditional Distribution of Test Statistics
-Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$.
-Then, we consider the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$ conditional on the event that the anomalous region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomalous region $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$, i.e.,
-$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (14)
+In this case, the $p$-value called naive $p$-value, and defined as
+$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\right).$ (13)
+If we can identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$, we can compute a valid $p$-value that control the false positive detection rate (i.e., the type I error rate).
+4 Selective inference for diffusion-based anomaly localization
+In this section, we introduce selective inference (SI) framework for testing the anomalous regions detected by a diffusion model and propose a method to perform valid hypotheses tests.
+4.1 Computing valid $p$-values via selective inference
+Complexity of sampling distribution
+As mentioned in §3, we need to identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$ to compute the $p$-values.
+If we ignore the fact that the anomalous region is identified by a diffusion model, the test statistic in (11) is simply a linear transformation of the Gaussian random vectors $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$, and hence itself follows a Gaussian distribution:
+$\begin{gathered}T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}),\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.\end{gathered}$ (14)
+However, as mentioned in §3, in reality the dependence of $\bm{\nu}$ on $\bm{x}$ via the diffusion model is more intricate, making the sampling distribution intractably complex.
+Consequently, obtaining this sampling distribution directly is challenging.
+Selective $p$-value via conditional sampling distribution
+Then, we consider the sampling distribution of the test statistic conditional on the event that the anomalous region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
+$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}}\right\}.$ (15)
 In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic.
 As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space.
-The nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
-$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}.$ (15)
-The $p$-value calculated from this conditional sampling distribution is called a selective $p$-value. Specifically, the selective $p$-value is defined as
-$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (16)
-where $\mathcal{X}$ is the conditional data space defined as
+The sufficient statistics of the nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
+$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}\bm{\nu}^{\top}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (16)
+where $I_{2n}$ is the identity matrix of size $2n$. By additional conditioning on the nuisance parameter, the selective $p$-value is defined as
+$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right).$ (17)
+The following theorem establishes that the selective $p$-value is a valid $p$-value for controlling the false positive detection rate for any significance level $\alpha\in(0,1)$.
+Theorem 1.
+The selective $p$-value in (17) is valid for controlling the false positive detection rate, i.e,
+$\mathbb{P}_{\mathrm{H}_{0}}\left(p_{\mathrm{selective}}\leq\alpha\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right)=\alpha,\;\forall\alpha\in(0,1).$ (18)
+Then, the selective $p$-value satisfies the following condition:
+$\mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha)=\alpha,\;\forall\alpha\in(0,1).$ (19)
+The proof of Theorem 1 is given in Appendix A.1.
+The following theorem tells that the selective $p$-value can be analytically derived from the conditional sampling distribution, which follows a truncated Gaussian distribution.
+Theorem 2.
+Consider the truncation intervals defined as
+$\mathcal{Z}=\left\{z\in\mathbb{R}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\},$ (20)
+where $X(z)$ are defined as
+$\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z,\;\;\text{where}\;\;\bm{a}=\mathcal{Q}_{\bm{x}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}},$ (21)
+and $\bm{a}_{1:n}$ and $\bm{b}_{1:n}$ denote the first $n$ elements of $\bm{a},\bm{b}\in\mathbb{R}^{2n}$ , respectively. Then, the selective $p$-value in (17) can be rewritten as
+$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}\right).$ (22)
+The conditional sampling distribution of the test statistic $T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))\mid\{Z\in\mathcal{Z}\}$ follows a truncated Gaussian distribution $\mathcal{TN}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu})$.
+The proof of the Theorem 2 is given in Appendix A.2.
+Once the truncation intervals $\mathcal{Z}$ are identified, computing the selective $p$-value in (22) becomes straightforward.
+Therefore, the remaining task is the identification of $\mathcal{Z}$.
+4.2 Identification of subinterval $\mathcal{Z}^{\mathrm{sub}}$ in $\mathcal{Z}$
+To tackle the identification of the truncation invervals $\mathcal{Z}$, we employ a divide-and-conquer strategy.
+Directly characterizing $\mathcal{Z}$ is challenging due to the complexity of the diffusion model’s computation.
+Our method decomposes the $n$-dimensional data space into a collection of polyhedra by imposing additional conditioning, a process we refer to as over-conditioning (OC) Duy and Takeuchi (2022).
+Each polyhedron in the $n$-dimensional space corresponds to an interval in the one-dimensional space $\mathcal{Z}$.
+Thus, we can examine these intervals sequentially in $\mathcal{Z}$ to determine whether they yield the same selected anomalous regions as observed.
+To this end, we need to identify a subinterval $\mathcal{Z}^{\mathrm{sub}}\subseteq\mathcal{Z}$.
+We show that the anomalous region from a diffusion model can be characterized by a piecewise-linear mapping, and that for each $z\in\mathbb{R}$, the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed analytically by solving a system of linear inequalities (see Appendix B for details).
+4.3 Identification of $\mathcal{Z}$
+Over-conditioning causes a reduction in power due to excessive conditioning.
+A technique called parametric programming is utilized to explore all intervals along the one-dimensional line, resulting in (20).
+The truncation intervals $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{sub}}$ as
+$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}}\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z).$ (23)
+An algorithm for computing the selective $p$-value is summarized in Algorithm 2.
+Figure 2 illustrates the example of conditional sampling distribution.
+It shows how the subintervals determine the conditional sampling distribution.
+Figure 2: Schematic illustration of the selective inference procedure for diffusion models.
+It shows how the image $\bm{X}(z)$ changes with $z$.
+The truncation intervals that yield the same anomalous region $\mathcal{M}_{\bm{X}(z)}$ as the observed anomalous region $\mathcal{M}_{\bm{x}}$ define the conditional sampling distribution.
+The $p_{\mathrm{selective}}$ denotes the proportion of probability mass within the truncation intervals.
+Algorithm 2 Selective $p$-value by Parametric Programming
+0:  $\bm{x},\bm{x}^{\mathrm{ref}}$
+1:  Initialize $\mathcal{Z}\leftarrow\emptyset$
+2:  Compute $\mathcal{M}_{\bm{x}}$, $\bm{a}$, $\bm{b}$, and $T(\bm{x},\bm{x}^{\mathrm{ref}})$ by (8), (21), (11)
+3:  Initialize $z$ to a sufficiently small value
+4:  while $z$ is not large enough do
+5:     Compute $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (34)
+6:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}$ then
+7:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$
+8:     end if
+9:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)}+\gamma$, where $0<\gamma\ll 1$
+10:  end while
+11:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}\right)$
+11:  $p_{\mathrm{selective}}$
+5 Experiments
+We compared our proposed method (proposed) on type I error rate and power with the following methods (see Appendix C for details of those methods):
+  • w/o-pp: An ablation study without the parametric programming technique described in §4.3.
+The $p$-value is computed using (22), with $\mathcal{Z}$ replaced by $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})$.
+  • naive: This method is conventionally used in practice. The $p$-value is computed as (13).
+  • bonferroni: This method uses the Bonferroni correction. Bonferroni correction is widely used for multiple testing correction.
+  • permutation: This method uses the permutation test. A permutation test is widely used for non-parametric hypothesis testing.
+The architecture of diffusion models used in all experiments is detailed in Appendix D. We executed all experiments on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
+5.1 Numerical experiments
+Experimental setup
+Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$.
+In the type I error rate experiments, we used only normal images.
+The synthetic dataset for normal images is generated to follow $\bm{X}\sim\mathcal{N}(\bm{0},\Sigma)$.
+We made 1,000 normal images for $n\in\left\{64,256,1024,4096\right\}$.
+In the power experiments, we used only abnormal images.
+We generated 1,000 abnormal images $\bm{X}\sim\mathcal{N}(\bm{\mu},\Sigma)$.
+The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen.
+The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$.
+In all experiments, we made the synthetic dataset for 1,000 reference images to follow $\bm{X}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma)$.
+The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$.
+All experiments were conducted under the significance level $\alpha=0.05$.
+The diffusion models were trained on the normal images from the synthetic dataset.
+The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings (see Appendix E for details).
+The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear.
+In all experiment, we generated normal-looking images through probabilistic sampling, $\eta$ was set to $1$.
+In addition, we conducted robustness experiments against non-Gaussian noise.
+The details of the robustness experiments are described in Appendix F.
+Results
+Figure 4 shows the comparison results of type I error rates.
+The proposed methods proposed and w/o-pp can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$.
+In contrast, naive and permutation cannot control the type I error rate.
+Figure 4 shows the comparison results of powers.
+Since naive and permutation cannot control the type I error rate, their powers are not considered.
+Among the methods that can control the type I error rate, the proposed has the highest power.
+The ablation study w/o-pp is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
+(a) Independence
+(b) Correlation
+Figure 3: Type I error rate comparison
+(a) Independence
+(b) Correlation
+Figure 4: Power comparison
+5.2 Real-world data experiments
+Experimental setup
+We conducted experiments using T2-FLAIR MRI brain scans from the Brain Tumor Segmentation (BraTS) 2023 dataset Karargyris et al.
+(2023); LaBella et al.
+(2023) and MVtec AD dataset Bergmann et al.
+(2019).
+The details of the experimental settings are described in Appendix G.
+Results
+Table 1: Type I error rate and power comparison on real-world datasets at the significance level of $\alpha(=0.05)$.
+The proposed and bonferroni methods control the type I error rate at or below $\alpha$ for each dataset, whereas naive fails to do so and is therefore unreliable.
+Each entry shows type I error rate (left) and power (right).
+Additionally, proposed achieves higher power than bonferroni.
+The Figures 9 and 10 show the results of the proposed and naive methods for the MRI images (BraTS) and MVTec AD dataset, respectively.
+Dataset naive bonferroni proposed
+Bottle 0.46/N.A 0.00/0.00 $\bm{0.04}$/$\bm{0.18}$
+Cable 0.88/N.A 0.00/0.00 $\bm{0.02}$/$\bm{0.40}$
+Grid 0.82/N.A 0.00/0.00 $\bm{0.06}$/$\bm{0.34}$
+Transistor 0.86/N.A 0.00/0.00 $\bm{0.08}$/$\bm{0.28}$
+BraTS (T2-FLAIR) 0.59/N.A 0.00/0.00 $\bm{0.05}$/$\bm{0.28}$
+Table 1 shows the comparison of the type I error rate and power.
+The naive cannot control the type I error rate, while the proposed and bonferroni can control the type I error rate below the significance level $\alpha$.
+The proposed has higher power than bonferroni.
+In addition, the examples of the results for each image are shown in Appendix H.
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.527$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
+(a) bottle (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$
+(b) Brain (Left: Normal, Right: Anomaly)
+Figure 5: Real-world examples of the naive and proposed methods.
+For each sample, the $p_{\mathrm{selective}}$ is high (true negative) for normal images and low (true positive) for abnormal images.
+In contrast, the $p_{\mathrm{naive}}$ remains low for all images, indicating that it fails to control the type I error rate.
+6 Conclusions, limitations and future works
+We introduced the DAL-Test, a novel statistical procedure for anomaly localizations identified by a diffusion model.
+With the proposed DAL-Test, the false positive detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by using a diffusion model.
+We demonstrated that the DAL-Test has higher power than the bonferroni correction, the only existing method for controlling the false positive detection rate.
+However, the growing the size of the diffusion model also leads to increased computational demands.
+In future work, we will focus on improving the computational efficiency of the DAL-Test.
+Acknowledgement
+This work was partially supported by MEXT KAKENHI (20H00601), JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JST AIP Acceleration Research (JPMJCR21U2), NEDO (JPNP18002, JPNP20006) and RIKEN Center for Advanced Intelligence Project.
+Appendix A Proofs
+A.1 Proof of Theorem 1
+Under the null hypothesis, the probability integral transform implies that
+$p_{\mathrm{selective}}\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\sim\mathrm{Uniform}(0,1),$ (24)
+and hence for any $\alpha\in(0,1)$,
+$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\right)=\alpha,\;\forall\alpha\in(0,1).$ (25)
+By marginalizing over the nuisance parameter $\mathcal{Q}_{\bm{x}}$, we have
+$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\right)$ (26)
+$=\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}})\ \mathbb{P}_{\mathrm{H}_{0}}(\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}})d\mathcal{Q}_{\bm{x}}$
 […]
-Due to the conditioning on the nuisance parameter $\mathcal{Q}_{\bm{X}}$, the conditional data space $\mathcal{X}$ can be rewritten as
-$\mathcal{X}=\left\{\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}\in\mathbb{R}^{2n}\middle|\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (18)
-where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$, and $\bm{c}_{1:n}$ represents a vector composed of the first $n$ elements of the vector $\bm{c}$.
-The vectors $\bm{a},\bm{b}\in\mathbb{R}^{2n}$ are defined as
-$\bm{a}=\mathcal{Q}_{\bm{X}_{\mathrm{obs}}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}},$ (19)
-and the region $\mathcal{Z}$ is defined as
-$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (20)
-Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z_{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}_{1:n}+\bm{b}_{1:n}Z$ and $\bm{X}_{\mathrm{obs}}=\bm{a}_{1:n}+\bm{b}_{1:n}z_{\mathrm{obs}}$.
-Then, the selective $p$-value in (16) is re-written as
-$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (21)
-Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}})$.
-Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution.
-Once the truncated region $\mathcal{Z}$ is identified, computing the selective $p$-value in (21) becomes straightforward.
-Therefore, the remaining task is the identification of $\mathcal{Z}$.
-4.2 Over-Conditioning
-To compute the truncated region $\mathcal{Z}$, we employ a divide and conquer approach.
-It is difficult to directly identify the truncated region $\mathcal{Z}$ due to the complexity of the computational algorithm of the diffusion model.
-The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022).
-It is easy to understand that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$.
-Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomalous region) as the observed one is selected.
-In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$.
-By exploiting this, we identify a over-conditioned interval $\mathcal{Z}^{\mathrm{oc}}\subset\mathcal{Z}$.
-Identification of $\mathcal{Z}^{\mathrm{oc}}$.
-Let us write a polyhedron $\mathcal{P}$ composed of piecewise-linear functions as
-$\mathcal{P}_{k}=\left\{\bm{\Delta}_{k}\bm{X}\leq\bm{\delta}_{k}\right\},k\in[K],$ (22)
-where $\bm{\Delta}_{k}$ and $\bm{\delta}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions of the $k$-th piecewise-linear function, respectively.
-Then, a piecewise-linear function $\mathcal{A}(\bm{X})$ is written in the following form:
-$\mathcal{A}(\bm{X})=\begin{cases}\bm{\Psi}_{1}\bm{X}+\bm{\psi}_{1}&\text{if }\bm{X}\in\mathcal{P}_{1},\\ \bm{\Psi}_{2}\bm{X}+\bm{\psi}_{2}&\text{if }\bm{X}\in\mathcal{P}_{2},\\ \hskip 28.45274pt\vdots\\ \bm{\Psi}_{K}\bm{X}+\bm{\psi}_{K}&\text{if }\bm{X}\in\mathcal{P}_{K},\end{cases}$ (23)
-where $\bm{\Psi}_{k}$ and $\bm{\psi}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions for the $k$-th polyhedron, respectively.
-Using the notation in (18), since the input image $\bm{X}(z)$ is restricted on a one-dimensional line, each component of the output of $\mathcal{A}$ is written as
-$\begin{gathered}\mathcal{A}_{i}(\bm{X}(z))=\begin{cases}\kappa^{\mathcal{A}_{i}}_{1}z+\rho^{\mathcal{A}_{i}}_{1}&\text{if }[L^{\mathcal{A}_{i}}_{1},U^{\mathcal{A}_{i}}_{1}],\\ \kappa^{\mathcal{A}_{i}}_{2}z+\rho^{\mathcal{A}_{i}}_{2}&\text{if }[L^{\mathcal{A}_{i}}_{2},U^{\mathcal{A}_{i}}_{2}],\\ \hskip 28.45274pt\vdots\\ \kappa^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}z+\rho^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}&\text{if }[L^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})},U^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}],\end{cases}\end{gathered}$ (24)
-where $K(\mathcal{A}_{i})$ is the number of linear pieces of the piecewise-linear function, and $\kappa^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ and $\rho^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ for $k\in[K(\mathcal{A}_{i})]$ are the coefficient and the constant of the $k$-th polyhedron, respectively.
-For each $i\in[n]$, there exists $k\in[K(\mathcal{A}_{i})]$ such that $z\in[L^{\mathcal{A}_{i}}_{k},U^{\mathcal{A}_{i}}_{k}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
-$\begin{gathered}\coloneqq\begin{cases}\left[\max\left(L^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right),U^{\mathcal{A}_{i}}_{k}\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}>0,\\ \left[L^{\mathcal{A}_{i}}_{k},\min\left(U^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right)\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}<0.\end{cases}\end{gathered}$ (25)
-We denote the over-conditioned interval as
-$\mathcal{Z}^{\text{oc}}(\bm{a}+\bm{b}z)=\bigcap_{i\in[n]}\left[L^{i}_{z},U^{i}_{z}\right].$ (26)
-Piecewise linearity of diffusion models.
-We now show that the diffusion model mapping $\mathcal{D}$ and then filtered reconstruction error $E$ can be expressed as a piecewise-linear function of $\bm{X}$.
+$=\alpha$
+Therefore, we have
+$\mathbb{P}_{\mathrm{H}_{0}}\left(p_{\mathrm{selective}}\leq\alpha\right)$ (27)
+$=\sum_{\mathcal{M}_{\bm{x}}\in 2^{[n]}}\mathbb{P}_{\mathrm{H}_{0}}(\mathcal{M}_{\bm{x}})\ \mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}})$
+$=\alpha\sum_{\mathcal{M}_{\bm{x}}\in 2^{[n]}}\mathbb{P}_{\mathrm{H}_{0}}(\mathcal{M}_{\bm{x}})$
+$=\alpha$
+A.2 Proof of Theorem 2
+The conditioning on $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}$ implies
+$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\Leftrightarrow\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}\bm{\nu}^{\top}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\Leftrightarrow\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z,$ (28)
+where $z=T(\bm{X},\bm{X}^{\mathrm{ref}})\in\mathbb{R}$. Hence,
+$\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}$ (29)
+=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z\right\}$
+=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\}$
+=$=\left\{\bm{a}+\bm{b}z\;\middle|\;z\in\mathcal{Z}\right\},$
+where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$. As a result
+$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\sim\mathcal{TN}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu})$ (30)
+Appendix B Calculating the subinterval $\mathcal{Z}^{\text{sub}}$ for diffusion models
+We now show that a reconstruction error $\mathcal{E}$ via diffusion models can be expressed as a piecewise-linear function of $\bm{X}$.
 To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below.
 It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$.
-To make the reverse process a piecewise-linear function, we employ a U-Net architecture composed of piecewise-linear components such as ReLU activation function and average pooling.
+To make the reverse process a piecewise-linear function, we employ a U-Net composed of piecewise-linear components such as ReLU activation functions and pooling layers.
 Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$.
@@
 As a result, the entire mapping ${\mathcal{D}}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$.
-Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=E(\bm{X}))$ is piecewise-linear.
-By exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed.
-4.3 Identification of $\mathcal{Z}$ by Parametric Programming
-Over-conditioning causes a reduction in power due to excessive conditioning.
-A technique called Parametric Programming is utilized to explore all intervals along the one-dimensional line, resulting in (20).
-The truncated region $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{oc}}$ as
-$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (27)
-The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them.
-In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z_{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z_{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$.
-An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
-Algorithm 2 Selective $p$-value Computation by Parametric Programming
-0:  $\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z_{\mathrm{obs}}:=T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})$
-1:  $\mathcal{Z}\leftarrow\emptyset$
-2:  Obtain $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ by (7)
-3:  Compute $\bm{a}$, $\bm{b}$ by (20)
-4:  $z\leftarrow z_{\mathrm{min}}$
-5:  while $z<z_{\mathrm{max}}$ do
-6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (26) for $z$
-7:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ then
-8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$
-9:     end if
-10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)}+\gamma$, where $\gamma$ is small positive number.
-11:  end while
-12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
-12:  $p_{\mathrm{selective}}$
-5 Experiments
-We compared our proposed methods (DMAD-test, DMAD-test-oc) with the other methods: naive method (naive), bonferroni correction (bonferroni), and permutation test (permutation) on type I error rate and power.
-The details of the methods for comparison are described in Appendix B.
-We executed the experiment on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
-5.1 Numerical Experiments
-Experimental setup.
-Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$.
-In the type I error rate experiments, we used only normal images.
-The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$.
-We made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$.
-In the power experiments, we used only abnormal images.
-We generated 1000 abnormal images $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{\mu},\Sigma)$.
-The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen.
-The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$.
-In all experiments, we made the synthetic dataset for 1000 reference images to follow $\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$.
-The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$.
-All experiments were conducted under the significance level $\alpha=0.05$.
-The details of the architecture of the diffusion model we used are shown in Appendix C.
-The diffusion models were trained on the normal images from the synthetic dataset.
-The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings.
-The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear.
-In all experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$.
-In addition, we conducted robustness experiments against non-Gaussian noise.
+Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=\mathcal{E}(\bm{X}))$ is piecewise-linear.
+Exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed.
+The following theorem tells that the subinterval $\mathcal{Z}^{\text{sub}}(\bm{a}+\bm{b}z)$ can be computed by solving a set of linear inequalities.
+Theorem 3.
+The piecewise-linear mapping $\mathcal{A}(\bm{X})$ can be expressed as a linear function of the input image $\bm{X}$ on each polyhedral region $\mathcal{P}_{k}$.
+$\forall\bm{X}\in\mathcal{P}^{(k)},\enskip\mathcal{A}(\bm{X})=\bm{\delta}^{(k)}+\bm{\Delta}^{(k)}\bm{X},$ (31)
+where $\bm{\delta}^{(k)}$ and $\bm{\Delta}^{(k)}$ for $k\in[K]$ are the constant vector and the coefficient matrix with appropriate dimensions for the $k$-th polyhedron, respectively.
+Using the notation in (22), since the input image $\bm{X}(z)$ is restricted on a one-dimensional line, each component of the output of $\mathcal{A}$ is written as
+$\forall z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}],\enskip\mathcal{A}_{i}(\bm{X}(z))=\kappa^{(k^{\prime})}_{i}+\rho^{(k^{\prime})}_{i}z,$ (32)
+where $\kappa^{(k^{\prime})}_{i}\in\mathbb{R}$ and $\rho^{(k^{\prime})}_{i}\in\mathbb{R}$ for $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ are the coefficient and the constant of the $k^{\prime}$-th interval $[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, and $K^{{}^{\prime}}_{i}$ is the number of linear pieces of $\mathcal{A}_{i}$.
+For each $i\in[n]$, there exists $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ such that $z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
+$(z)\begin{gathered}\coloneqq\begin{cases}\left[\max\left(L^{(k^{\prime})}_{i},(\lambda-\rho^{(k^{\prime})}_{i})/\kappa^{(k^{\prime})}_{i}\right),U^{(k^{\prime})}_{i}\right]&\text{if }\kappa^{(k^{\prime})}_{i}>0,\\ \left[L^{(k^{\prime})}_{i},\min\left(U^{(k^{\prime})}_{i},(\lambda-\rho^{(k^{\prime})}_{i})/\kappa^{(k^{\prime})}_{i}\right)\right]&\text{if }\kappa^{(k^{\prime})}_{i}<0.\end{cases}{}\end{gathered}$ (33)
+By applying the above theorem, we denote the subinterval as
+$\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)=\bigcap_{i\in[n]}\left[L_{i}(z),U_{i}(z)\right].$ (34)
+Appendix C Comparison methods for numerical experiments
+We compared our proposed method with the following methods:
+  • proposed: The proposed method uses the parametric programming.
+  • w/o-pp: The proposed method use over-conditioning (without parametric programming). The $p$-value is calculated as
+    $p_{\mathrm{ablation}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|>|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})\right)$ (35)
+  • naive: The naive method. This method uses a conventional $z$-test without any conditioning in (13).
+  • bonferroni: To control the type I error rate, this method applies the Bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
+    $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
+  • permutation: This method uses a permutation test with the steps outlined below:
+    • Calculate the observed test statistic $z^{\mathrm{obs}}$ by applying the observed image $\bm{x}$ to the diffusion model.
+    • For each $i=1,\ldots,B$, compute the test statistic $z^{(i)}$ by applying the permuted image $\bm{X}^{(i)}$ to the diffusion model, where $B$ represents the total number of permutations, set to 1,000 in our experiments.
+      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z^{\mathrm{obs}}|\},$
+      where $\bm{1}\{\cdot\}$ denotes the indicator function.
+This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
+Appendix D Architecture of the U-Net
+Figure 6 shows the architecture of the U-Net used in our experiments.
+The U-Net has three skip connections, and the Encoder and Decoder blocks.
+For image sizes $n\in\{64,256,1024,4096\}$, the corresponding spatial dimensions of images are $(1,d,d)$ where $d\in\{8,16,32,64\}$.
+Figure 6: The architecture of the U-Net
+Appendix E Accelerated reverse processes
+Methods for accelerating the reverse process have been proposed in DDPM, DDIM Song et al.
+(2022).
+When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$.
+In this case, equations (3) and (5) can be rewritten as
+$\mathbf{x}_{\tau_{i-1}}=\sqrt{\alpha_{\tau_{i-1}}}\left(\frac{\mathbf{x}_{\tau_{i}}-\sqrt{1-\alpha_{\tau_{i}}}\cdot\epsilon^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})}{\sqrt{\alpha_{\tau_{i}}}}\right)+\sqrt{1-\alpha_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\cdot\epsilon_{\theta}^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})+\sigma_{\tau_{i}}\epsilon_{\tau_{i}},$ (36)
+where
+$\sigma_{\tau_{i}}=\eta\sqrt{(1-\alpha_{\tau_{i-1}})/(1-\alpha_{\tau_{i}})}\sqrt{1-\alpha_{\tau_{i}}/\alpha_{\tau_{i-1}}}.$ (37)
+Therefore, piecewise-linearity is preserved, making the proposed method DMAD-test applicable.
+Appendix F Robustness of the proposed method
 To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$.
@@
 Figure 7 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$.
-We run $1000$ trials for each distribution family and each 1-Wasserstein distance to calculate the type I error rate.
 The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$.
-Figure 2: Non-Gaussian distributions with $d=0.04$
-Results.
-Figure 4 shows type I error rate comparison.
-The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$.
-In contrast, naive and permutation cannot control the type I error rate.
-Figure 4 shows power comparison.
-Since naive and permutation cannot control the type I error rate, their powers are not considered.
-Among the methods that can control the type I error rate, the proposed method has the highest power.
-DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
 Figure 8 shows the results of the robustness experiments.
 DMAD-test maintains good performance on the type I error rate for all the considered distribution families.
-(a) Independence
-(b) Correlation
-Figure 3: Type I Error Rate comparison
-(a) Independence
-(b) Correlation
-Figure 4: Power comparison
+Figure 7: Non-Gaussian distributions with $d=0.04$
 (a) SND
@@
 Figure 8: Type I Error Rate for Non-Gaussian distribution families
-5.2 Real Data Experiments
-We examined the medical image experiment using the dataset extracted from the dataset in Buda et al.
+Appendix G Experimental settings for the MVTec AD dataset
+G.1 Experimental settings of Brain Trumor Segmentation 2023 dataset
+We evaluate our method on T2-FLAIR MRI brain scans from the Brain Trumor Segmentation 2023 dataset Karargyris et al.
+(2023); LaBella et al.
 (2023).
-This dataset includes brain images without tumors as normal images (941 images) and brain images with tumors as abnormal images (939 images).
-A random selection of 100 images from the normal images was made to estimate the mean and variance.
-The remaining images were standardized using the estimated mean and variance, and 329 normal images were extracted for training the diffusion model.
+T2-FLAIR MRI which comprises 934 non-skull-stripped 3D scans with dimensions of $240\!\times\!240\!\times\!155$.
+From these scans, we extracted 2D $240\!\times\!240$ axial slices at axis $95$, resized them to $64\!\times\!64$ pixels, and categorized them based on the anomaly annotations into $532$ normal images (without tumors) and $402$ abnormal images (with tumors).
+For each scan, we estimated the mean and variance from pixel values excluding both the non-brain regions and tumor regions identified in the ground truth, followed by standardization.
+We randomly selected $312$ normal images for model training.
 The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings.
-We set the threshold $\lambda=0.9$ and the kernel size of the averaging filter to $3$.
-The results of the DMAD-test and naive are shown in the Figure 6.
-The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors.
-The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for abnormal images (True Positive), indicating that the results are desirable.
-On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for abnormal images but also for normal images (False Positive), indicating the invalidness of the naive test.
-This result indicates that the DMAD-test detected anomalous regions as statistically significant while avoiding misidentification of normal image as anomaly.
-$p_{\rm naive}=0.019$, $p_{\rm selective}=0.643$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.379$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.290$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.822$ Refer to caption $p_{\rm naive}=0.018$, $p_{\rm selective}=0.373$ Refer to caption $p_{\rm naive}=0.024$, $p_{\rm selective}=0.932$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.225$
-(a) Brain images without tumors
-$p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.002$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.021$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.028$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.027$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.016$
-(b) Brain images with tumors
-Figure 6: An example of the results for applying the proposed DMAD-test and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to brain images.
-6 Conclusions
-In this study, we proposed a statistical test for anomalous regions in medical images detected by a diffusion model.
-With the proposed DMAD-test, the false detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by a diffusion model.
-We believe this study marks a step toward bridging the gap between generative AI and rigorous statistical inference in medical imaging analysis.
-Acknowledgement
-This work was partially supported by MEXT KAKENHI (20H00601), JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JST AIP Acceleration Research (JPMJCR21U2), NEDO (JPNP18002, JPNP20006) and RIKEN Center for Advanced Intelligence Project.
-Appendix A Accelerated Reverse Processes
-Methods for accelerating the reverse process have been proposed in DDPM, DDIM (Song et al., 2022).
-When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$.
-In this case, equations (2) and (4) can be rewritten as
-$\mathbf{x}_{\tau_{i-1}}=\sqrt{\alpha_{\tau_{i-1}}}\left(\frac{\mathbf{x}_{\tau_{i}}-\sqrt{1-\alpha_{\tau_{i}}}\cdot\epsilon^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})}{\sqrt{\alpha_{\tau_{i}}}}\right)+\sqrt{1-\alpha_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\cdot\epsilon_{\theta}^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})+\sigma_{\tau_{i}}\epsilon_{\tau_{i}},$ (28)
-where
-$\sigma_{\tau_{i}}=\eta\sqrt{(1-\alpha_{\tau_{i-1}})/(1-\alpha_{\tau_{i}})}\sqrt{1-\alpha_{\tau_{i}}/\alpha_{\tau_{i-1}}}.$ (29)
-Therefore, piecewise-linearity is preserved, making the proposed method DMAD-test applicable.
-Appendix B Comparison Methods for Numerical Experiments
-We compared our proposed method with the following methods:
-  • DMAD-test: The proposed method uses the parametric programming.
-  • DMAD-test-oc: The proposed method uses over-conditioning.
-  • naive: The naive method. This method uses a conventional $z$-test without any conditioning. The naive $p$-value is calculated as
-    $p_{\mathrm{naive}}=\mathbb{P}_{\mathbb{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\right).$
-  • bonferroni: To control the type I error rate, this method applies the bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
-    $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
-  • permutation: This method uses a permutation test with the steps outlined below:
-    • Calculate the observed test statistic $z_{\mathrm{obs}}$ by applying the observated image $\bm{X}_{\mathrm{obs}}$ to the diffusion model.
-    • For each $i=1,\ldots,B$, compute the test statistic $z^{(i)}$ by applying the permuted image $\bm{X}^{(i)}$ to the diffusion model, where $B$ represents the total number of permutations, set to 1,000 in our experiments.
-    • The permutation $p$-value is then determined as
-      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z_{\mathrm{obs}}|\},$
-      where $\bm{1}\{\cdot\}$ denotes the indicator function.
-This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
-Appendix C Architecture of the U-Net
-Figure 7 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks.
-Figure 7: The architecture of the U-Net
+We set the threshold $\lambda=0.6$ and the kernel size of the averaging filter to $3$.
+Note that, when testing images of the MRI brain scans, the non-brain regions are not treated as anomalous regions $\mathcal{M}_{\bm{X}}$.
+G.2 Experimental settings of Brain Trumor Segmentation 2023 dataset
+We evaluate our method on the MVTec AD dataset Bergmann et al.
+(2019), which consists of $15$ object categories.
+Each category provides a training set of normal images and a test set containing both normal and abnormal images, with image resolutions ranging from $900\times 900$ to $1024\times 1024$ pixels.
+For our experiments, we select six categories (bottle, cable, grid) and resize all images to $128\times 128$ pixels.
+For the type I error rate experiments, we randomly select $50$ normal images from each category, and for the power experiments, we select $50$ abnormal images per category.
+The diffusion model is trained on the remaining normal images with $T=1000$ total diffusion steps, of which the first $T^{\prime}=300$ steps are used for reconstruction using $4$ sampling steps.
+We apply an averaging filter with a kernel size of $3$.
+We set the anomaly threshold $\lambda$ to $1.0$ for bottle and to $1.2$ for cable and grid.
+To demonstrate in the power experiments, we compute the intersection between the anomalous region detected by the diffusion model and the anomaly annotation.
+Since the images in the MVTec AD dataset are RGB, we redefine the image data as $\tilde{\bm{X}}\in\mathbb{R}^{hw\times 3}$, where $h$ and $w$ denote the image height and width.
+We then vectorize each image by
+$\bm{X}$ $=\mathrm{vec}(\tilde{\bm{X}})$ (38)
+$=(X_{1,1},X_{1,2},X_{1,3},\,X_{2,1},X_{2,2},X_{2,3},\dots,X_{hw,1},X_{hw,2},X_{hw,3})^{\top}\in\mathbb{R}^{n},$
+Similarly, we define the reference image $\tilde{\bm{X}}^{\mathrm{ref}}\in\mathbb{R}^{hw\times 3}$ as the average of the training images, and vectorize it in the same way as above.
+$\bm{X}^{\mathrm{ref}}$ $=\mathrm{vec}(\tilde{\bm{X}}^{\mathrm{ref}})$ (39)
+$=(\tilde{X}_{1,1},\tilde{X}_{1,2},\tilde{X}_{1,3},\,\tilde{X}_{2,1},\tilde{X}_{2,2},\tilde{X}_{2,3},\dots,\tilde{X}^{\mathrm{ref}}_{hw,1},\tilde{X}^{\mathrm{ref}}_{hw,2},\tilde{X}^{\mathrm{ref}}_{hw,3})^{\top}\in\mathbb{R}^{n}$
+where $n=3hw$, and accordingly redefine the test statistic in (11) as
+$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\sum_{j\in[3]}\tilde{X}_{i,j}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\sum_{j\in[3]}\tilde{X}^{\mathrm{ref}}_{i,j}$ (40)
+With this setup, the our method can be applied in the same way as in §3.
+Appendix H Experimental results for the real-world datasets
+In this section, we show the experimental results for the real-world datasets. We applied the proposed proposed to the MRI images and MVTec AD dataset and compared with the naive method.
+$p_{\rm naive}:0.031$, $p_{\rm selective}:0.266$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.766$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.360$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.646$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.753$ Refer to caption $p_{\rm naive}:0.011$, $p_{\rm selective}:0.768$
+MRI images without tumors
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.026$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.012$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.007$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.002$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
+MRI images with tumors
+Figure 9: An example of the results for applying the proposed proposed and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to MRI images.
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.907$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.010$
+(a) Bottle (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.567$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
+(b) Cable (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.028$, $p_{\rm selective}:0.789$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
+(c) Grid (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.769$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
+(d) Transistor (Left: Normal, Right: Anomaly)
+Figure 10: An example of applying the proposed method and the naive method to six categories of the MVTec AD dataset (Bottle, Cable, Grid, Transistor).
+For each category, the left figure shows a normal image and the right figure shows an anomalous image.
+The proposed $p_{\rm selective}$ remains high for normal samples (True Negatives) and low for anomalous samples (True Positives), demonstrating accurate control of the false detection rate.
+In contrast, the $p_{\rm naive}$ yields low $p$-values across both normal and anomalous images, indicating an inflated false positive rate and invalidity of the naive method.

### v5

@@
-Quantifying Statistical Significance in Diffusion-Based Anomaly Localization via Selective Inference
+Statistical Test for Diffusion-Based Anomaly Localization via Selective Inference
 Abstract
-Anomaly localization in images—identifying regions that deviate from expected patterns—is vital in applications such as medical diagnosis and industrial inspection.
+Anomaly localization in images—identifying regions that deviate from normal patterns—is vital in applications such as medical diagnosis and industrial inspection.
 A recent trend is the use of image generation models in anomaly localization, where these models generate normal-looking counterparts of anomalous images, thereby allowing flexible and adaptive anomaly localization.
@@
 1 Introduction
-Anomaly localization using image generation models, particularly diffusion models, has shown great promise across diverse domains such as medical diagnosis and industrial inspection Li et al.
-(2023); Iqbal et al.
-(2023); Lu et al.
-(2023); Zhang et al.
-(2023); Fontanella et al.
-(2024); Tebbe and Tayyub (2024); Sheng et al.
-(2024).
+Anomaly localization using image generation models, particularly diffusion models, has shown great promise across diverse domains such as medical diagnosis and industrial inspection (Li et al., 2023; Iqbal et al., 2023; Lu et al., 2023; Zhang et al., 2023; Fontanella et al., 2024; Tebbe and Tayyub, 2024; Sheng et al., 2024).
 These models reconstruct a normal-looking version of an input image, and differences between the input and the reconstruction highlight potential anomalies.
-Compared to traditional methods, generative approaches are highly suitable for settings where annotations for anomalous regions are unavailable.
-Moreover, generative approaches can flexibly handle heterogeneity by adapting to individual images–e.g., patient-specific characteristics in medical diagnosis and product-specific traits in industrial inspection.
-Among various image generation models, diffusion models, in particular, offer high fidelity and stability, outperforming other methods in image quality and anomaly localization.
-While generative approaches offer powerful and flexible capabilities for anomaly localization, a major concern is that the inherent uncertainty and bias in generative models can affect localization performance Fithian et al.
-(2014); Taylor and Tibshirani (2015); Lee et al.
-(2016); Duy and Takeuchi (2022); Miwa et al.
-(2023); Shiraishi et al.
-(2024).
-These models are trained on specific datasets composed of normal images, and the quality of the generated normal-looking images depends heavily on the distribution of the dataset and how well the model has learned the underlying distribution.
+Compared to traditional anomaly localization methods, generative approaches are highly suitable for settings where annotations for anomalous regions are unavailable.
+Moreover, generative approaches can flexibly handle heterogeneity by adapting to individual images—e.g., patient-specific characteristics in medical diagnosis and product-specific traits in industrial inspection.
+Among various image generation models, diffusion models, in particular, offer high fidelity and stability, often outperforming other methods in image quality and anomaly localization.
+While generative approaches offer powerful and flexible capabilities for anomaly localization, a major concern is that the inherent uncertainty and bias in generative models can affect localization performance (Fithian et al., 2014; Taylor and Tibshirani, 2015; Lee et al., 2016; Duy and Takeuchi, 2022; Miwa et al., 2023; Shiraishi et al., 2024).
+These models are trained on specific datasets composed of normal images, and the quality of the generated normal-looking images depends heavily on the dataset distribution and how accurately the model has learned it.
 As a result, uncertainties or biases in the dataset or training process can cause incorrect reconstructions, leading to inaccurate localizations and misidentification of anomalies.
 Such risks are especially critical in high-stakes domains such as medical diagnosis and industrial inspection, where even minor errors can have serious consequences.
-Therefore, it is essential to incorporate rigorous uncertainty quantification framework and statistical safeguards to ensure reliable deployment in critical applications.
+Therefore, it is essential to incorporate a rigorous uncertainty quantification framework and statistical safeguards to ensure reliable deployment in critical applications.
 To address this issue, we propose a statistical testing framework based on Selective Inference (SI) to assess the statistical significance of detected anomalies.
-SI has recently emerged as a promising approach for conducting statistical inference on hypotheses that are selected based on observed data.
+SI has emerged as a promising approach for conducting statistical inference on hypotheses that are selected based on observed data.
 In this framework, inference is performed using the conditional distribution given the selection event, thereby accounting for the uncertainty and bias associated with the hypothesis selection.
@@
 This allows us to quantify the statistical significance of detected anomalies as a valid $p$-value, providing a rigorous estimate of the false positive rate and offering a principled metric for reliability.
-In this paper, as a proof of concept for the proposed statistical testing framework, we focus on a standard denoising diffusion probabilistic model (DDPM) Ho et al.
-(2020); Song et al.
-(2022) among various diffusion-based anomaly detection (AD) methods , and its applications to medical diagnostics and industrial inspections.
+In this paper, as a proof of concept for the proposed statistical testing framework, we focus on a standard denoising diffusion probabilistic model (DDPM) (Ho et al., 2020; Song et al., 2022) among various diffusion-based anomaly detection (AD) methods, and demonstrate its applications to medical diagnostics and industrial inspections.
 However, the proposed framework is readily generalizable to a broader range of diffusion model architectures and is applicable to semi-supervised AD problems in other domains.
 Related works
-Diffusion models have been effectively utilized in anomaly localization problems Pinaya et al.
-(2022); Fontanella et al.
-(2024); Wyatt et al.
-(2022); Mousakhan et al.
-(2023).
-In this context, the DDPM is commonly used Ho et al.
-(2020); Song et al.
-(2022).
+Diffusion models have been effectively utilized in anomaly localization problems (Pinaya et al., 2022; Fontanella et al., 2024; Wyatt et al., 2022; Mousakhan et al., 2023).
+In this context, the DDPM is commonly used (Ho et al., 2020; Song et al., 2022).
 During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise.
@@
 The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions.
-Other types of generative AI has also been used for anomalous region detection task Baur et al.
-(2021); Chen and Konukoglu (2018); Chow et al.
-(2020); Jana et al.
-(2022).
-SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm Lee and Taylor (2014); Lee et al.
-(2016); Tibshirani et al.
-(2016), and then extended to more complex feature selection methods Yang et al.
-(2016); Suzumura et al.
-(2017); Hyun et al.
-(2018); Rügamer and Greven (2020); Das et al.
-(2021).
-Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks Chen and Bien (2020); Tsukurimichi et al.
-(2021); Tanizaki et al.
-(2020); Duy et al.
-(2022); Le Duy et al.
-(2024); Lee et al.
-(2015); Gao et al.
-(2022); Duy et al.
-(2020); Jewell et al.
-(2022).
-The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data.
-To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event.
-To the best of our knowledge, SI was applied to statistical inferences on several deep learning models Duy et al.
-(2022); Miwa et al.
-(2023); Shiraishi et al.
-(2024); Miwa et al.
-(2024), but none of them works on image generation by diffusion models.
-Our contributions
-The main contributions of our study are summarized as follows.
-[^1] We note that our contribution is not the development of a new diffusion-based anomaly localization algorithm, but rather the introduction of a rigorous statistical testing framework designed to quantify the statistical reliability of anomalous regions identified by diffusion-based AD methods.
-  • We propose a novel statistical testing framework to assess the significance of anomaly localization results derived from diffusion model-based methods, offering a principled basis for evaluating the reliability of detected anomalies.
+Other types of generative models have also been used for the anomaly localization task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
+SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021).
+Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022).
+The fundamental idea of SI is to perform inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data.
+To conduct SI, it is necessary to derive the sampling distribution of the test statistic conditional on the hypothesis selection event.
+To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of these works addresses image generation by diffusion models.
+Contributions
+The main contributions of our study are summarized as follows [^1] We note that our contribution is not the development of a new diffusion-based anomaly localization algorithm, but rather the introduction of a rigorous statistical testing framework designed to quantify the statistical reliability of anomalous regions identified by diffusion-based AD methods..
+  • We propose a novel statistical testing framework to assess the significance of anomaly localization results obtained from diffusion model-based methods, offering a principled basis for evaluating the reliability of detected anomalies.
   • We implement the SI framework for diffusion models by deriving the sampling distribution conditional on the selection event induced by the diffusion model, which requires developing non-trivial computational techniques tailored to the generative sampling process.
@@
 The results highlight the robustness and practical utility of our method.
-The implementation code for reproducing all experimental results is provided at https://github.com/tkatsuoka/DAL-test
-Figure 1: Schematic illustration of anomaly localization on a brain MRI image dataset using a diffusion model and the proposed DAL-Test.
+The implementation code for reproducing all experimental results is available at https://github.com/tkatsuoka/DAL-test
+Figure 1: Schematic illustration of anomaly localization using a diffusion model and the proposed DAL-Test.
 When a test image—potentially containing an anomalous region—is fed into a trained diffusion model, its normal-looking version is generated through the forward and reverse processes.
@@
 An anomaly score is then computed based on the difference between the reconstructed image and the input test image, and the spatial distribution of this score is analyzed to localize anomalies.
-By applying thresholding to the score map, anomalous regions can be clearly identified.
-In this study, for the purpose of proof of concept, we adopt standard denoising diffusion models as our choice of diffusion model Ho et al.
-(2020); Song et al.
-(2022).
-The following outlines the image reconstruction process of a trained DDPM.
-Given a test image which possibly contain anomalous regions, a denoising diffusion model is used to generate the corresponding normal image.
-The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
+By applying thresholding to the score map, anomalous regions can be identified.
+In this study, for the purpose of proof of concept, we adopt standard denoising diffusion models as our choice of diffusion model (Ho et al., 2020; Song et al., 2022).
+We outline the image reconstruction process of a trained DDPM below.
+Given a test image that possibly contains anomalous regions, a denoising diffusion model is used to generate the corresponding normal image.
+The reconstruction consists of two processes called forward process (or diffusion process) and reverse process.
 In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$.
@@
 We consider the distribution of the original and noisy test images, which is denoted by $q(\bm{x})$, and approximate the distribution by a parametric model $p_{\theta}(\bm{x})$ with $\theta$ being the parameters.
-Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
+Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2}<\dots<\beta_{T}<1$, the forward process is written as
 $q(\bm{x}_{1:T}|\bm{x}_{0}):=\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1}),\;\;\text{where}\;\;q(\bm{x}_{t}|\bm{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}I).$ (1)
@@
 where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
-In the reverse process, a parametric model in the form of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\mu_{\theta}(\bm{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\bm{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$.
+In the reverse process, a parametric model in the form of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\mu_{\theta}(\bm{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\bm{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$.
 Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$.
-In DDPM Ho et al.
-(2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\bm{x}_{t})-\epsilon_{t}||^{2}_{2}$.
+In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $\epsilon^{(t)}_{\theta}(\bm{x}_{t})-\epsilon_{t}||^{2}_{2}$.
 Based on (2), given a noisy image $\bm{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\bm{x}_{t-1}$ is obtained as
@@
 In practice, the reverse process starts from $\bm{x}_{T^{\prime}}$ with $T^{\prime}<T$.
-Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contains individual information of the original input image.
+Namely, we reconstruct the normal-looking version of the input image not from a completely noisy one, but from one that still contains individual information of the original input image.
 The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image.
-Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions Ho et al.
-(2020); Mousakhan et al.
-(2023).
+Conversely, the larger $T^{\prime}$ results in the retention of only large-scale features, thereby converting more of the anomalous regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023).
+The overall reconstruction process of the diffusion model is summarized in Algorithm 1.
 Algorithm 1 Reconstruction Process
-0:  Input image $\bm{x}$
-1:  $\bm{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\bm{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
-2:  for $t=T^{\prime},\dots,1$ do
-3:     $f^{(t)}_{\theta}(\bm{x}_{t})\leftarrow(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}}$
-4:     $\bm{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t})+\sigma_{t}\epsilon_{t}$
-5:  end for
-5:  Reconstructed image $\bm{x}_{0}$
-3 Statistical test for diffusion-based anomalous localization
+1:Input image $\bm{x}$
+[…]
+3:for $t=T^{\prime},\dots,1$ do
+4:  $f^{(t)}_{\theta}(\bm{x}_{t})\leftarrow(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}}$
+5:  $\bm{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t})+\sigma_{t}\epsilon_{t}$
+6:end for
+7:Reconstructed image $\bm{x}_{0}$
+3 Statistical test for diffusion-based anomaly localization
 In this section, we formulate the statistical test for detecting anomalous regions by a trained diffusion model.
@@
 First, in the training phase, the diffusion model is trained only on normal images.
-Then, in the test phase, we feed a test image which might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$.
-By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image.
+Then, in the test phase, we feed a test image that might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$.
+By appropriately selecting $T^{\prime}$, we can generate a normal-looking image that retains individual characteristics of the test input image.
 If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image.
-On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions.
+On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to reconstruct these regions, as it has been trained primarily on normal regions.
 Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
@@
 We emphasize that the noise component $\bm{\varepsilon}$ should not be confused with the noise $\epsilon$ used in the forward process.
-Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint.
-On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and their covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI.
-Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
+No structural assumption is imposed on the signal $\bm{\mu}$.
+On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and its covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI.
+Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector,
 $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\varepsilon},\;\;\;\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (6)
 where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix.
-In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image vector is denoted as $\bm{x}$.
+In the following, we use $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image vector is denoted as $\bm{x}$.
 Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}\colon\mathbb{R}^{n}\ni\bm{X}\to\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$.
@@
 When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise.
-In this study, we simply used an averaging filter.
+In this study, we use an averaging filter.
 Let us denote the averaging filter as $\mathcal{F}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$.
@@
 $\mathcal{E}\colon\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (7)
-where absolute value is taken pixel-wise.
+where the absolute value is taken pixel-wise.
 Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $\mathcal{E}_{i}(\bm{X})$ for each pixel $i\in[n]$.
@@
 $\begin{gathered}\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\text{with}\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),\end{gathered}$ (9)
-where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image and the $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component.
+where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image, $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component, and $\bm{X}^{\mathrm{ref}}$ is assumed to be independent of $\bm{X}$.
 Then, we consider the following null and alternative hypotheses:
-$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i}\\ ~{}~{}~{}\text{v.s.}~{}~{}~{}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i},\end{gathered}$ (10)
-where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the observed anomalous region $\mathcal{M}_{\bm{x}}$, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different.
+$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i}\\ ~~~\text{vs.}~~~\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i},\end{gathered}$ (10)
+where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference image in the observed anomalous region $\mathcal{M}_{\bm{x}}$, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different.
 A reasonable test statistic for the statistical test in (10) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
@@
 $\bm{\nu}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{x}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{x}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (12)
-where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
-In this case, the $p$-value called naive $p$-value, and defined as
+where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 0 otherwise.
+In this case, the $p$-value is called the naive $p$-value, and is defined as
 $p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\right).$ (13)
-If we can identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$, we can compute a valid $p$-value that control the false positive detection rate (i.e., the type I error rate).
+If we can identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$, we can compute a valid $p$-value that controls the false positive detection rate (i.e., the type I error rate).
 4 Selective inference for diffusion-based anomaly localization
-In this section, we introduce selective inference (SI) framework for testing the anomalous regions detected by a diffusion model and propose a method to perform valid hypotheses tests.
+In this section, we introduce the SI framework for testing the anomalous regions detected by a diffusion model and propose a method to perform valid hypothesis tests.
 4.1 Computing valid $p$-values via selective inference
@@
 As mentioned in §3, we need to identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$ to compute the $p$-values.
-If we ignore the fact that the anomalous region is identified by a diffusion model, the test statistic in (11) is simply a linear transformation of the Gaussian random vectors $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$, and hence itself follows a Gaussian distribution:
-$\begin{gathered}T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}),\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.\end{gathered}$ (14)
+If we ignore the fact that the anomalous region is identified by a diffusion model, the test statistic in (11) is simply a linear transformation of the Gaussian random vectors $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$, and hence itself follows a Gaussian distribution under the null hypothesis $\mathrm{H}_{0}$:
+[…]
+where $\bm{X}$ and $\bm{X}^{\mathrm{ref}}$ are assumed to be independent.
 However, as mentioned in §3, in reality the dependence of $\bm{\nu}$ on $\bm{x}$ via the diffusion model is more intricate, making the sampling distribution intractably complex.
@@
 As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space.
-The sufficient statistics of the nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
+The sufficient statistic of the nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
 $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}\bm{\nu}^{\top}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}$, (16)
@@
 The proof of Theorem 1 is given in Appendix A.1.
-The following theorem tells that the selective $p$-value can be analytically derived from the conditional sampling distribution, which follows a truncated Gaussian distribution.
+The following theorem states that the selective $p$-value can be analytically derived from the conditional sampling distribution, which follows a truncated Gaussian distribution.
 Theorem 2.
@@
 $\mathcal{Z}=\left\{z\in\mathbb{R}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\},$ (20)
-where $X(z)$ are defined as
+where $\bm{X}(z)$ is defined as
 $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z,\;\;\text{where}\;\;\bm{a}=\mathcal{Q}_{\bm{x}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}}{\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu}},$ (21)
@@
 The conditional sampling distribution of the test statistic $T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))\mid\{Z\in\mathcal{Z}\}$ follows a truncated Gaussian distribution $\mathcal{TN}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu})$.
-The proof of the Theorem 2 is given in Appendix A.2.
+The proof of Theorem 2 is given in Appendix A.2.
 Once the truncation intervals $\mathcal{Z}$ are identified, computing the selective $p$-value in (22) becomes straightforward.
 Therefore, the remaining task is the identification of $\mathcal{Z}$.
-4.2 Identification of subinterval $\mathcal{Z}^{\mathrm{sub}}$ in $\mathcal{Z}$
-To tackle the identification of the truncation invervals $\mathcal{Z}$, we employ a divide-and-conquer strategy.
+4.2 Identification of truncation intervals $\mathcal{Z}$
+Identification of subinterval $\mathcal{Z}^{\mathrm{sub}}$
+To tackle the identification of the truncation intervals $\mathcal{Z}$, we employ a divide-and-conquer strategy.
 Directly characterizing $\mathcal{Z}$ is challenging due to the complexity of the diffusion model’s computation.
-Our method decomposes the $n$-dimensional data space into a collection of polyhedra by imposing additional conditioning, a process we refer to as over-conditioning (OC) Duy and Takeuchi (2022).
+Our method decomposes the $n$-dimensional data space into a collection of polyhedra by imposing additional conditioning, a process we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022).
 Each polyhedron in the $n$-dimensional space corresponds to an interval in the one-dimensional space $\mathcal{Z}$.
@@
 To this end, we need to identify a subinterval $\mathcal{Z}^{\mathrm{sub}}\subseteq\mathcal{Z}$.
-We show that the anomalous region from a diffusion model can be characterized by a piecewise-linear mapping, and that for each $z\in\mathbb{R}$, the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed analytically by solving a system of linear inequalities (see Appendix B for details).
-4.3 Identification of $\mathcal{Z}$
+We show that the anomalous region from a diffusion model can be characterized by a piecewise-linear mapping.
+Exploiting this property, the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed analytically for each $z\in\mathbb{R}$ by solving a system of linear inequalities (see Appendix B).
+Identification of $\mathcal{Z}$ via parametric programming
 Over-conditioning causes a reduction in power due to excessive conditioning.
@@
 $\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}}\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z).$ (23)
-An algorithm for computing the selective $p$-value is summarized in Algorithm 2.
-Figure 2 illustrates the example of conditional sampling distribution.
-It shows how the subintervals determine the conditional sampling distribution.
+Figure 2 illustrates the conditional sampling distribution determined by these truncation intervals. An algorithm for computing the selective $p$-value is summarized in Algorithm 2.
+Algorithm 2 Selective $p$-value by Parametric Programming
+[…]
+2:Initialize $\mathcal{Z}\leftarrow\emptyset$
+3:Compute $\mathcal{M}_{\bm{x}}$, $\bm{a}$, $\bm{b}$, and $T(\bm{x},\bm{x}^{\mathrm{ref}})$ by (8), (11), (21)
+4:Initialize $z$ to a sufficiently small value
+5:while $z$ is not large enough do
+6:  Compute $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (34)
+7:  if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}$ then
+8:   $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$
+9:  end if
+10:  $z\leftarrow\max{\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)}+\gamma$, where $0<\gamma\ll 1$
+11:end while
+[…]
+[…]
 Figure 2: Schematic illustration of the selective inference procedure for diffusion models.
@@
 The $p_{\mathrm{selective}}$ denotes the proportion of probability mass within the truncation intervals.
-Algorithm 2 Selective $p$-value by Parametric Programming
-0:  $\bm{x},\bm{x}^{\mathrm{ref}}$
-1:  Initialize $\mathcal{Z}\leftarrow\emptyset$
-2:  Compute $\mathcal{M}_{\bm{x}}$, $\bm{a}$, $\bm{b}$, and $T(\bm{x},\bm{x}^{\mathrm{ref}})$ by (8), (21), (11)
-3:  Initialize $z$ to a sufficiently small value
-4:  while $z$ is not large enough do
-5:     Compute $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (34)
-6:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}$ then
-7:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$
-8:     end if
-9:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)}+\gamma$, where $0<\gamma\ll 1$
-10:  end while
-11:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|\geq|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}\right)$
-11:  $p_{\mathrm{selective}}$
 5 Experiments
 We compared our proposed method (proposed) on type I error rate and power with the following methods (see Appendix C for details of those methods):
-  • w/o-pp: An ablation study without the parametric programming technique described in §4.3.
+  • w/o-$p$: An ablation study without the parametric programming technique described in §4.2.
 The $p$-value is computed using (22), with $\mathcal{Z}$ replaced by $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})$.
@@
   • permutation: This method uses the permutation test. A permutation test is widely used for non-parametric hypothesis testing.
-The architecture of diffusion models used in all experiments is detailed in Appendix D. We executed all experiments on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
+The architecture of diffusion models used in all experiments is detailed in Appendix D. We executed all experiments on AMD EPYC 9474F processor, 48-core 3.63.6GHz CPU and 768768GB memory.
 5.1 Numerical experiments
@@
 The synthetic dataset for normal images is generated to follow $\bm{X}\sim\mathcal{N}(\bm{0},\Sigma)$.
-We made 1,000 normal images for $n\in\left\{64,256,1024,4096\right\}$.
+We generated 1,000 normal images for $n\in\left\{64,256,1024,4096\right\}$.
 In the power experiments, we used only abnormal images.
@@
 The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$.
-In all experiments, we made the synthetic dataset for 1,000 reference images to follow $\bm{X}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma)$.
+In all experiments, we generated the synthetic dataset for 1,000 reference images to follow $\bm{X}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma)$.
 The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$.
@@
 The diffusion models were trained on the normal images from the synthetic dataset.
-The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings (see Appendix E for details).
+The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was performed using $5$-step sampling (see Appendix E for details).
 The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear.
-In all experiment, we generated normal-looking images through probabilistic sampling, $\eta$ was set to $1$.
-In addition, we conducted robustness experiments against non-Gaussian noise.
-The details of the robustness experiments are described in Appendix F.
+In all experiments, we generated normal-looking images through probabilistic sampling, with $\eta$ set to $1$.
 Results
@@
 Experimental setup
-We conducted experiments using T2-FLAIR MRI brain scans from the Brain Tumor Segmentation (BraTS) 2023 dataset Karargyris et al.
-(2023); LaBella et al.
-(2023) and MVtec AD dataset Bergmann et al.
-(2019).
-The details of the experimental settings are described in Appendix G.
+We conducted experiments using T2-FLAIR MRI brain scans from the Brain Tumor Segmentation (BraTS) 2023 dataset (Karargyris et al., 2023; LaBella et al., 2023) and MVTec AD dataset (Bergmann et al., 2019).
+The details of the experimental settings are described in Appendix F.
 Results
@@
 The proposed and bonferroni methods control the type I error rate at or below $\alpha$ for each dataset, whereas naive fails to do so and is therefore unreliable.
-Each entry shows type I error rate (left) and power (right).
 Additionally, proposed achieves higher power than bonferroni.
-The Figures 9 and 10 show the results of the proposed and naive methods for the MRI images (BraTS) and MVTec AD dataset, respectively.
-Dataset naive bonferroni proposed
-Bottle 0.46/N.A 0.00/0.00 $\bm{0.04}$/$\bm{0.18}$
-Cable 0.88/N.A 0.00/0.00 $\bm{0.02}$/$\bm{0.40}$
-Grid 0.82/N.A 0.00/0.00 $\bm{0.06}$/$\bm{0.34}$
-Transistor 0.86/N.A 0.00/0.00 $\bm{0.08}$/$\bm{0.28}$
-BraTS (T2-FLAIR) 0.59/N.A 0.00/0.00 $\bm{0.05}$/$\bm{0.28}$
+Figure 5 shows qualitative results of the proposed and naive methods on the MVTec AD dataset and the BraTS dataset.
+naive bonferroni proposed
+Dataset Type I Error Power Type I Error Power Type I Error Power
+Bottle 0.46 N.A 0.00 0.00 $\bm{0.04}$ $\bm{0.18}$
+Cable 0.88 N.A 0.00 0.00 $\bm{0.02}$ $\bm{0.40}$
+Grid 0.82 N.A 0.00 0.00 $\bm{0.06}$ $\bm{0.34}$
+Transistor 0.86 N.A 0.00 0.00 $\bm{0.08}$ $\bm{0.28}$
+BraTS (T2-FLAIR) 0.59 N.A 0.00 0.00 $\bm{0.05}$ $\bm{0.28}$
 Table 1 shows the comparison of the type I error rate and power.
@@
 The proposed has higher power than bonferroni.
-In addition, the examples of the results for each image are shown in Appendix H.
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.527$
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
-(a) bottle (Left: Normal, Right: Anomaly)
 $p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$
 $p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$
-(b) Brain (Left: Normal, Right: Anomaly)
-Figure 5: Real-world examples of the naive and proposed methods.
-For each sample, the $p_{\mathrm{selective}}$ is high (true negative) for normal images and low (true positive) for abnormal images.
-In contrast, the $p_{\mathrm{naive}}$ remains low for all images, indicating that it fails to control the type I error rate.
-6 Conclusions, limitations and future works
-We introduced the DAL-Test, a novel statistical procedure for anomaly localizations identified by a diffusion model.
-With the proposed DAL-Test, the false positive detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by using a diffusion model.
+(a) Brain MRI (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.907$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.010$
+(b) Bottle (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.567$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
+(c) Cable (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.028$, $p_{\rm selective}:0.789$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
+(d) Grid (Left: Normal, Right: Anomaly)
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.769$
+$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
+(e) Transistor (Left: Normal, Right: Anomaly)
+Figure 5: Results of the proposed and naive methods on the BraTS dataset (one example) and the MVTec AD dataset (four categories).
+For each dataset, the left image is a normal sample and the right image is an anomalous sample.
+6 Conclusions
+We introduced the DAL-Test, a novel statistical procedure for anomaly localization identified by a diffusion model.
+With the proposed DAL-Test, the false positive detection rate can be controlled at the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by using a diffusion model.
 We demonstrated that the DAL-Test has higher power than the Bonferroni correction, the only existing method for controlling the false positive detection rate.
-However, the growing the size of the diffusion model also leads to increased computational demands.
-In future work, we will focus on improving the computational efficiency of the DAL-Test.
 Acknowledgement
@@
 Under the null hypothesis, the probability integral transform implies that
-$p_{\mathrm{selective}}\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\sim\mathrm{Uniform}(0,1),$ (24)
+$p_{\mathrm{selective}}\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\sim\mathrm{Uniform}(0,1),$ (24)
 and hence for any $\alpha\in(0,1)$,
-$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}}\right\}\right)=\alpha,\;\forall\alpha\in(0,1).$ (25)
+$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\right)=\alpha,\;\forall\alpha\in(0,1).$ (25)
 By marginalizing over the nuisance parameter $\mathcal{Q}_{\bm{x}}$, we have
-$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\right)$ (26)
-$=\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}})\ \mathbb{P}_{\mathrm{H}_{0}}(\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{x}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}})d\mathcal{Q}_{\bm{x}}$
-[…]
+$\mathbb{P}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\,\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right)$ (26)
+$=\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}\!\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\,\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right)$
+$\qquad\qquad\cdot\,\mathbb{P}_{\mathrm{H}_{0}}\!\left(\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}}\right)d\mathcal{Q}_{\bm{x}}$
+$=\alpha\int_{\mathbb{R}^{n}}\mathbb{P}_{\mathrm{H}_{0}}\!\left(\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}}\right)d\mathcal{Q}_{\bm{x}}$
 $=\alpha$
@@
 $\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}$ (29)
-=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z\right\}$
-=$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\}$
-=$=\left\{\bm{a}+\bm{b}z\;\middle|\;z\in\mathcal{Z}\right\},$
-where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$. As a result
+$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}=\bm{a}+\bm{b}z\right\}$
+$=\left\{\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}\;\middle|\;\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{x}}\right\}$
+$=\left\{\bm{a}+\bm{b}z\;\middle|\;z\in\mathcal{Z}\right\},$
+where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$. As a result,
 $T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{x}},\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\mathcal{Q}_{\bm{x},\bm{x}^{\mathrm{ref}}}\right\}\sim\mathcal{TN}(0,\bm{\nu}^{\top}\tilde{\Sigma}\bm{\nu})$ (30)
 Appendix B Calculating the subinterval $\mathcal{Z}^{\text{sub}}$ for diffusion models
-We now show that a reconstruction error $\mathcal{E}$ via diffusion models can be expressed as a piecewise-linear function of $\bm{X}$.
+We show that a reconstruction error $\mathcal{E}$ via diffusion models can be expressed as a piecewise-linear function of $\bm{X}$.
 To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below.
@@
 Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=\mathcal{E}(\bm{X}))$ is piecewise-linear.
-Exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed.
-The following theorem tells that the subinterval $\mathcal{Z}^{\text{sub}}(\bm{a}+\bm{b}z)$ can be computed by solving a set of linear inequalities.
+Exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\mathrm{sub}}$ can be computed.
+The following theorem states that the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed by solving a set of linear inequalities.
 Theorem 3.
@@
 $\forall z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}],\enskip\mathcal{A}_{i}(\bm{X}(z))=\kappa^{(k^{\prime})}_{i}+\rho^{(k^{\prime})}_{i}z,$ (32)
-where $\kappa^{(k^{\prime})}_{i}\in\mathbb{R}$ and $\rho^{(k^{\prime})}_{i}\in\mathbb{R}$ for $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ are the coefficient and the constant of the $k^{\prime}$-th interval $[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, and $K^{{}^{\prime}}_{i}$ is the number of linear pieces of $\mathcal{A}_{i}$.
+where $\kappa^{(k^{\prime})}_{i}\in\mathbb{R}$ and $\rho^{(k^{\prime})}_{i}\in\mathbb{R}$ for $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ are the coefficient and the constant of the $k^{\prime}$-th interval $L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}$], and $K^{{}^{\prime}}_{i}$ is the number of linear pieces of $\mathcal{A}_{i}$.
 For each $i\in[n]$, there exists $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ such that $z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}$], then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
@@
 We compared our proposed method with the following methods:
-  • proposed: The proposed method uses the parametric programming.
-  • w/o-pp: The proposed method use over-conditioning (without parametric programming). The $p$-value is calculated as
+  • proposed: The proposed method uses parametric programming.
+  • w/o-pp: The proposed method uses over-conditioning (without parametric programming). The $p$-value is calculated as
     $p_{\mathrm{ablation}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X}(Z),\bm{X}^{\mathrm{ref}}(Z))|>|T(\bm{x},\bm{x}^{\mathrm{ref}})|\;\middle|\;Z\in\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z^{\mathrm{obs}})\right)$ (35)
   • naive: The naive method. This method uses a conventional $z$-test without any conditioning in (13).
-  • bonferroni: To control the type I error rate, this method applies the Bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
+  • bonferroni: To control the type I error rate, this method applies the Bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as
     $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
@@
       where $\bm{1}\{\cdot\}$ denotes the indicator function.
-This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
 Appendix D Architecture of the U-Net
@@
 Appendix E Accelerated reverse processes
-Methods for accelerating the reverse process have been proposed in DDPM, DDIM Song et al.
-(2022).
+Methods for accelerating the reverse process have been proposed in DDPM, DDIM (Song et al., 2022).
 When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$.
@@
 $\sigma_{\tau_{i}}=\eta\sqrt{(1-\alpha_{\tau_{i-1}})/(1-\alpha_{\tau_{i}})}\sqrt{1-\alpha_{\tau_{i}}/\alpha_{\tau_{i-1}}}.$ (37)
-Therefore, piecewise-linearity is preserved, making the proposed method DMAD-test applicable.
-Appendix F Robustness of the proposed method
-To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$.
-We considered the following non-Gaussian distributions with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
-  • Skew normal distribution family (SND).
-  • Exponentially modified gaussian distribution family (EMG).
-  • Generalized normal distribution family (GND) with a shape parameter $\beta$. This distribution family can be steeper than the normal distribution (i.e., $\beta<2$).
-  • Student’s $t$-distribution family ($t$-distribution).
-Note that these distributions are standardized in the experiments.
-Figure 7 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$.
-The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$.
-Figure 8 shows the results of the robustness experiments.
-DMAD-test maintains good performance on the type I error rate for all the considered distribution families.
-Figure 7: Non-Gaussian distributions with $d=0.04$
-(a) SND
-(b) EMG
-(c) GND
-(d) $t$-distribution
-Figure 8: Type I Error Rate for Non-Gaussian distribution families
-Appendix G Experimental settings for the MVTec AD dataset
-G.1 Experimental settings of Brain Trumor Segmentation 2023 dataset
-We evaluate our method on T2-FLAIR MRI brain scans from the Brain Trumor Segmentation 2023 dataset Karargyris et al.
-(2023); LaBella et al.
-(2023).
-T2-FLAIR MRI which comprises 934 non-skull-stripped 3D scans with dimensions of $240\!\times\!240\!\times\!155$.
+Therefore, piecewise-linearity is preserved, making the proposed method DAL-Test applicable.
+Appendix F Experimental settings for the real-world datasets
+F.1 Experimental settings of Brain Tumor Segmentation 2023 dataset
+We evaluate our method on T2-FLAIR MRI brain scans from the Brain Tumor Segmentation 2023 dataset (Karargyris et al., 2023; LaBella et al., 2023).
+T2-FLAIR MRI comprises 934 non-skull-stripped 3D scans with dimensions of $240\!\times\!240\!\times\!155$.
 From these scans, we extracted 2D $240\!\times\!240$ axial slices at axis $95$, resized them to $64\!\times\!64$ pixels, and categorized them based on the anomaly annotations into $532$ normal images (without tumors) and $402$ abnormal images (with tumors).
@@
 Note that, when testing images of the MRI brain scans, the non-brain regions are not treated as anomalous regions $\mathcal{M}_{\bm{X}}$.
-G.2 Experimental settings of Brain Trumor Segmentation 2023 dataset
-We evaluate our method on the MVTec AD dataset Bergmann et al.
-(2019), which consists of $15$ object categories.
+F.2 Experimental settings of MVTec AD dataset
+We evaluate our method on the MVTec AD dataset (Bergmann et al., 2019), which consists of $15$ object categories.
 Each category provides a training set of normal images and a test set containing both normal and abnormal images, with image resolutions ranging from $900\times 900$ to $1024\times 1024$ pixels.
-For our experiments, we select six categories (bottle, cable, grid) and resize all images to $128\times 128$ pixels.
+For our experiments, we select four categories (bottle, cable, grid, transistor) and resize all images to $128\times 128$ pixels.
 For the type I error rate experiments, we randomly select $50$ normal images from each category, and for the power experiments, we select $50$ abnormal images per category.
@@
 We set the anomaly threshold $\lambda$ to $1.0$ for bottle and to $1.2$ for cable and grid.
-To demonstrate in the power experiments, we compute the intersection between the anomalous region detected by the diffusion model and the anomaly annotation.
+In the power experiments, we compute the intersection between the anomalous region detected by the diffusion model and the anomaly annotation.
 Since the images in the MVTec AD dataset are RGB, we redefine the image data as $\tilde{\bm{X}}\in\mathbb{R}^{hw\times 3}$, where $h$ and $w$ denote the image height and width.
 We then vectorize each image by
-$\bm{X}$ $=\mathrm{vec}(\tilde{\bm{X}})$ (38)
-$=(X_{1,1},X_{1,2},X_{1,3},\,X_{2,1},X_{2,2},X_{2,3},\dots,X_{hw,1},X_{hw,2},X_{hw,3})^{\top}\in\mathbb{R}^{n},$
+$\bm{X}=\mathrm{vec}(\tilde{\bm{X}})=(X_{1,1},X_{1,2},X_{1,3},\,X_{2,1},X_{2,2},X_{2,3},\dots,X_{hw,1},X_{hw,2},X_{hw,3})^{\top}\in\mathbb{R}^{n},$ (38)
 Similarly, we define the reference image $\tilde{\bm{X}}^{\mathrm{ref}}\in\mathbb{R}^{hw\times 3}$ as the average of the training images, and vectorize it in the same way as above.
-$\bm{X}^{\mathrm{ref}}$ $=\mathrm{vec}(\tilde{\bm{X}}^{\mathrm{ref}})$ (39)
-$=(\tilde{X}_{1,1},\tilde{X}_{1,2},\tilde{X}_{1,3},\,\tilde{X}_{2,1},\tilde{X}_{2,2},\tilde{X}_{2,3},\dots,\tilde{X}^{\mathrm{ref}}_{hw,1},\tilde{X}^{\mathrm{ref}}_{hw,2},\tilde{X}^{\mathrm{ref}}_{hw,3})^{\top}\in\mathbb{R}^{n}$
+$\bm{X}^{\mathrm{ref}}=\mathrm{vec}(\tilde{\bm{X}}^{\mathrm{ref}})=(\tilde{X}_{1,1},\tilde{X}_{1,2},\tilde{X}_{1,3},\,\tilde{X}_{2,1},\tilde{X}_{2,2},\tilde{X}_{2,3},\dots,\tilde{X}^{\mathrm{ref}}_{hw,1},\tilde{X}^{\mathrm{ref}}_{hw,2},\tilde{X}^{\mathrm{ref}}_{hw,3})^{\top}\in\mathbb{R}^{n}$ (39)
 where $n=3hw$, and accordingly redefine the test statistic in (11) as
 $T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\sum_{j\in[3]}\tilde{X}_{i,j}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\sum_{j\in[3]}\tilde{X}^{\mathrm{ref}}_{i,j}$ (40)
-With this setup, the our method can be applied in the same way as in §3.
-Appendix H Experimental results for the real-world datasets
-In this section, we show the experimental results for the real-world datasets. We applied the proposed proposed to the MRI images and MVTec AD dataset and compared with the naive method.
-$p_{\rm naive}:0.031$, $p_{\rm selective}:0.266$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.766$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.360$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.646$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.753$ Refer to caption $p_{\rm naive}:0.011$, $p_{\rm selective}:0.768$
-MRI images without tumors
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.026$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.012$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.007$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.002$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
-MRI images with tumors
-Figure 9: An example of the results for applying the proposed proposed and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to MRI images.
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.907$
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.010$
-(a) Bottle (Left: Normal, Right: Anomaly)
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.567$
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
-(b) Cable (Left: Normal, Right: Anomaly)
-$p_{\rm naive}:0.028$, $p_{\rm selective}:0.789$
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
-(c) Grid (Left: Normal, Right: Anomaly)
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.769$
-$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
-(d) Transistor (Left: Normal, Right: Anomaly)
-Figure 10: An example of applying the proposed method and the naive method to six categories of the MVTec AD dataset (Bottle, Cable, Grid, Transistor).
-For each category, the left figure shows a normal image and the right figure shows an anomalous image.
-The proposed $p_{\rm selective}$ remains high for normal samples (True Negatives) and low for anomalous samples (True Positives), demonstrating accurate control of the false detection rate.
-In contrast, the $p_{\rm naive}$ yields low $p$-values across both normal and anomalous images, indicating an inflated false positive rate and invalidity of the naive method.
+With this setup, our method can be applied in the same way as in §3.